### v1

# Eagle: Exploring The Design Space for Multimodal LLMs with Mixture of Encoders

Abstract
The ability to accurately interpret complex visual information is a crucial topic of multimodal large language models (MLLMs). Recent work indicates that enhanced visual perception significantly reduces hallucinations and improves performance on resolution-sensitive tasks, such as optical character recognition and document analysis. A number of recent MLLMs achieve this goal using a mixture of vision encoders. Despite their success, there is a lack of systematic comparisons and detailed ablation studies addressing critical aspects, such as expert selection and the integration of multiple vision experts. This study provides an extensive exploration of the design space for MLLMs using a mixture of vision encoders and resolutions. Our findings reveal several underlying principles common to various existing strategies, leading to a streamlined yet effective design approach. We discover that simply concatenating visual tokens from a set of complementary vision encoders is as effective as more complex mixing architectures or strategies. We additionally introduce Pre-Alignment to bridge the gap between vision-focused encoders and language tokens, enhancing model coherence. The resulting family of MLLMs, Eagle, surpasses other leading open-source models on major MLLM benchmarks.
Figure 1: Overview of Eagle. Eagle is a family of multimodal large language models (MLLMs) with a mixture of vision encoders. Left: comparisons between Eagle and existing competitive MLLMs with Vicuna-7B [15] and Llama3-8B [3], with Eagle achieving favorable results on all 12 benchmarks. Middle: an evolutionary road map of the design space and advanced training recipes leading to consistent and significant improvements. Right: highlights and core features of Eagle.
1 Introduction
The success of large language models (LLMs) has triggered significant interest in enabling their visual perception capability, such that they could see, understand, and reason in the real world. At the center of these multimodal large language models (MLLMs) [23] is a typical design where images are converted into a series of visual tokens by the vision encoders and appended with the text embeddings. CLIP [63] is often chosen as the vision encoder since its visual representation is align with the text space by pre-training on image-text pairs. Depending on the architectures, training recipes and the way how vision tokens are injected into the language model, there exist various notable families of MLLMs such as Flamingo [4], BLIP [44, 43, 16], PaLI [11], PaLM-E [18] and LLaVA [54, 52]. Most of these works keep relatively low input resolutions due to the limits on pre-trained vision encoders and LLM sequence length.
Recent studies [47, 53] show that stronger vision encoder design is important for mitigating MLLM hallucinations [50, 77] and improving resolution-sensitive tasks like optical character recognition (OCR). A constellation of works thus focuses on enhancing the capability of the vision encoder. For example, scaling up the pre-training data and parameters of vision encoder [13] or dividing images into low-resolution patches [53, 67]. However, these approaches usually introduce large training resources. An efficient yet powerful strategy is to mixing visual encoders pre-trained with different tasks and input resolutions, either fusing higher resolution encoders with the CLIP encoder [59, 45], sequentially appending features from different encoders [20, 49], or adopting more complex fusion and routing strategies to make the best of different encoders [37, 88]. Such a “mixture-of-vision-experts” mode is shown to be effective. However, a detailed study of its design space with rigorous ablation is still lacking, which motivates us to revisiting this area. Questions such as which vision encoder combination to choose, how to fuse different experts together, and how to adjust training strategies with more vision encoders remain to be answered.
To address the above questions, our work systematically investigates the mixture-of-vision-encoders design space for improved MLLM perception. As shown in Fig. 1, our exploration of the design space consists of the following steps: 1) Benchmarking various vision encoders and searching recipes for higher resolution adaptation; 2) “Apples to apples” comparison between vision encoder fusion strategies; 3) Progressive identification of the optimal combination of multiple vision encoders; 4) Improved vision expert pre-alignment and data mixture. Our study covers the performance of vision encoders pre-trained on different tasks and resolutions (e.g., vision-language alignment [31, 14, 63, 66], self-supervised learning [62], detection [22, 21], segmentation [36], and OCR [38]). We use a round-robin approach to incorporate additional vision experts. Starting with the basic CLIP [63] encoder, we add one additional expert at a time with the best improvement in each round.
Our work is not the first one to leverage multiple vision encoders in MLLM. However, the systematic study leads to several interesting new findings under this setting:
  • Unlocking the vision encoders during MLLM training matters. This is in sharp contrast to the LLaVA [54, 52] family and many works that consider multiple vision encoders or teachers [49, 55, 20, 33, 64, 37], where freezing the vision encoders has been a common choice.
  • Some recently proposed fusion strategies [59, 45] do not show significant advantages despite their advanced designs. Instead, we find that straightforward channel concatenation stands out as a simple yet competitive fusion strategy, offering the best efficiency and performance.
  • Incorporating additional vision experts leads to consistent gain, making it a promising path to systematic enhance MLLM perception besides scaling up single encoders. The improvement is particularly pronounced when vision encoders are unlocked.
  • We propose a pre-alignment stage where non-text-aligned vision experts are individually fine-tuned with a frozen LLM before trained together. This stage is found to enhance the MLLM performance significantly under the mixture-of-vision-encoder design.
We finally conclude our findings into a family of MLLMs termed Eagle. Eagle is evaluated on a series of benchmarks, including visual question answering, OCR/document related tasks, and benchmarks tailored for MLLMs. Our model attains state-of-the-art performance across different benchmarks and demonstrates obvious advantages on OCR and document understanding tasks. Using the same pre-train and supervised fine-tuning data from Cambrian-1 [73] - a concurrent family of vision-centric MLLMs sharing similar design spirits, Eagle models overall achieve better performance. We hope that the Eagle can provide a highly performant and easy-to-reproduce MLLM solution to the community.
2 Design space exploration
In this section, we show how to utilize the advantages of different vision encoders via step-by-step investigations, yielding the Eagle model family. Unlike previous methods focusing on new fusion strategies or architectures among vision encodes, our goal is to identify a set of minimalistic design to fuse different vision encoders supported with detailed ablations, removing any unnecessary parts. As shown in Fig. 2, we start from extending the basic CLIP encoder [63] to a set of vision experts with different architectures, pre-training tasks and resolutions. With these experts, we then compare different fusion architectures and methods, and study how to optimize the pre-training stratgies given more encoders. We also give a detailed analysis in how to select the vision encoders to be integrated. Finally, we put all the findings together and further extend to multiple expert vision encoders with different resolutions and domain knowledge.
2.1 Base setup
We adopt LLaVA’s [54] model architecture as the basis, which consists of a large language model, a vision encoder, and a projection layer. The projection layer projects the visual embedding from the vision encoder into the text embedding space. We mainly focus on the vision encoder design.
Figure 2: Overview of the Eagle exploration pipeline. We investigate the MLLM design space with multiple vision encoders, aiming to identify optimized design choices and improve the MLLM perception. We include various types of vision experts pre-trained on different tasks and resolutions. We then systematically compare a series of popular fusion strategies under controlled settings, including Sequence Append, Channel Concatenation, LLaVA-HR [59], Mini-Gemini [45], and Deformable Attention [86]. Finally, the optimal fusion design is further extended to multiple vision encoders to form a strong MLLM perception.
Base training data. We use the same pre-training data as LLaVA-$1.51.5$ [52], which consists of 595595595595k image text pairs. For the supervised fine-tuning stage, we collect data from a series of tasks and convert them into multimodal conversations, including: LLaVA-1.5 [52], Laion-GPT4V [1], ShareGPT-4V [8], DocVQA [61], synDog-EN [35], ChartQA [60], DVQA [32], and AI2D [34], resulting in 934k samples.
Implementation details. We first pre-train the model with image-text pairs for one epoch with a batch size of 256, where the whole model is frozen and only the projector layer is updated. In the second stage, we further fine-tune the model on the supervised fine-tuning data for one epoch with a batch size of 128. For this exploration, we employ Vicuna-7B as the underlying language model. The learning rates are set to 1e-3 for the first stage and 2e-5 for the second stage, respectively.
2.2 Stronger CLIP encoder
We start our exploration with the CLIP model since it has become the main choice for many MLLMs. While CLIP models are known to benefit multimodal tasks, their drawbacks have also been well-studied. For instance, many existing MLLMs [52] tend to use the pre-trained CLIP resolutions (such as $224\times 224224 × 224$ or $336\times 336336 × 336$) as their input resolutions. In these cases, the encoders often fail to capture fine-grained details that are important for resolution-sensitive tasks like OCR and document understanding [47].
To handle increased input resolution, a common practice is to use tiling where input images are divided into tiles and encoded separately [53, 47]. Another simpler method is to directly scale up the input resolution and interpolate the position embeddings of the vision transformer model if needed [9, 10, 6]. We compare these two approaches with frozen/unfrozen vision encoders under different resolutions, with the results shown in Table 1. Our findings can be summarized as follows:
  • Unfreezing the CLIP encoder leads to significant improvement when interpolating to a higher MLLM input resolution that is different from the CLIP pre-training resolution. There is also no performance degradation when resolutions remain the same.
  • When the CLIP encoder is frozen, directly adapting it to a higher MLLM input resolution considerably hurts the performance.
  • Among the compared strategies, directly interpolating to $448\times 448448 × 448$ with an unfrozen CLIP encoder is shown to be both effective and efficient in terms of performance and cost.
  • The best CLIP encoder gets close to InternVL in performance despite the significantly smaller model size (300M vs. 6B) and less pre-training data.
It is worth noting that CLIP-448 allows us to match the setting with LLaVA-HR [59] and InternVL [13] where the CLIP encoders are similarly adapted to take $448\times 448448 × 448$ input and output $10241024$ patch tokens [^1] The latest InternVL 1.0 and InternVL 1.2 models also adopt a pixel shuffle module. We remove it in our experiment to ensure full alignment. See https://github.com/OpenGVLab/InternVL for details.. In the following investigation, we stick to this simple strategy of scaling up the input resolution and unlocking the vision encoder during training.
Table 1: Comparison of different high-resolution adaption methods. Direct interpolation is more efficient and effective. “Res” denotes input image resolution. “#Token/s” denotes the inference speed of the whole framework. “#Token (V)” denotes the number of visual feature tokens fed into LLM. “OCR” denotes OCRBench [57]. “Avg” denotes average. Note that CLIP-448 gets very close to InternVL in terms of performance, despite the significantly smaller model size (300M vs. 6B).
Method Unfreeze  Res #Token/s #Token (V)  GQA  MME MMMU  OCR  SQA POPE  Avg
Original 336 47.2 576 62.46 1488 36.1 369 72.79 86.77 615
Original 336 47.2 576 63.68 1496 36.6 369 72.22 85.9 617
Interpolate 448 47.9 1024 62.45 1484 34.4 285 72.27 86.54 598
Interpolate 448 47.9 1024 64.62 1531 35.7 496 72.77 87.88 645
Interpolate 672 44.5 2304 64.51 1492 36.2 492 72.86 87.49 641
Tiled-input 672 44.8 2304 63.59 1455 34.8 445 71.94 87.83 625
InternVL [13] 448 46.2 1024 65.12 1521 36.2 527 72.46 87.13 649
Table 2: Detailed setting of the pre-trained vision experts.
Expert Category Resolution Token Number Post-Processing Variant Hyper-links
CLIP Image-Text Matching 448 1024 None VIT-L
ConvNeXt Image Classification 1024 1024 None ConvNeXt-XXL
EVA-02 Object Detection 1024 1024 Resize EVA-02-Large
Pix2Struct Text Recognition 1024 1024 Resize Pix2Struct-02-Large
DINOv2 Self-supervised 448 1024 None dinov2_vitl14_reg
SAM Image Segmentation 1024 1024 Pixel-unshuffle SAM-Large
2.3 Vision experts
To better establish the foundation for multi-vision expert fusion, we extend the toolbox with vision experts pre-trained on different tasks and resolutions, and verify our findings on high-resolution adaptation with these experts. This also helps us identify the distinct advantages of different experts. We collect a set of vision experts, including: (1) Vision-Language Alignment: CLIP [63] and ConvNeXt [76] from OpenCLIP [31, 66]. (2) Object-Centric: EVA-02 [22, 21] pre-trained on detection datasets. (3) OCR: Pix2Struct [38].(4) Segmentation: SAM [36]. (5) Self-supervised: DINOv2 [62]. The detailed input resolution and checkpoint of each vision encoder can be found in Table 2. We resize the output 2D feature maps of each vision encoder using bilinear interpolation and pixel shuffle [68] to ensure that the visual token number equals $10241024$.
The results in Table 3 reveal that MLLMs with these task-specific vision encoders achieve optimal performance in their pre-training domains. EVA-02 excels in the object hallucination evaluation benchmark POPE [46] and general visual question answering benchmark GQA [30]. CLIP and ConvNeXt perform well across all benchmarks, benefiting from their training on large-scale image-text pairs using contrastive loss. Conversely, while Pix2Struct excels in text recognition, it shows limited capability in object recognition and general VQA tasks, like POPE and GQA. DINOv2 and SAM, pre-trained with self-supervised learning and semantic segmentation respectively, struggle with text recognition tasks. Note that the conclusion made in Sec. 2.2 aligns with the results here. Both ConvNeXt and EVA-02 show significant performance gains when the vision encoders are updated during training.
Table 3: Comparison between different vision experts as the MLLM encoders. “Res” denotes resolution. “OCR” denotes OCRBench [57].
Category Vision Tower Unfreeze    Res  GQA  MME MMMU  OCR  SQA POPE TextVQA
VL Alignment ConvNeXt 1024 63.71 1473 34.2 402 71.92 87.42 66.95
1024 63.71 1433 37.0 527 72.01 87.91 71.66
Segmentation SAM 1024 54.87 1193 34.8 32 70.34 83.31 45.04
1024 60.05 1305 35.9 38 70.64 86.62 45.71
Object Detection EVA-02 1024 63.54 1422 35.8 58 70.86 88.50 47.86
1024 64.93 1474 35.7 387 71.78 88.96 59.79
Text Recognition Pix2Struct 1024 52.47 1217 35.7 443 69.61 77.54 56.56
1024 54.05 1220 34.3 425 70.29 79.21 57.05
Self-supervised DINOv2 448 60.81 1269 34.0 39 70.76 86.42 45.71
448 64.06 1371 34.2 37 71.35 87.29 46.77
2.4 Fusion strategy
Existing MLLM frameworks have proposed various fusion strategies to ensemble the encoders in Table 3, with the hope that their domain-specific strengths can be leveraged. In all cases, improvements in MLLM performance have been reported with the fusion of vision encoders. However, the roles of the fusion strategies as part of their MLLM architecture innovations, have not been decoupled and clearly studied under an “apples to apples” comparison. It is thus not entirely clear how much improvement is from the fusion strategies themselves versus the improved representations from various encoders.
We notice that existing popular fusion strategies, despite their variations in designs, can be broadly represented by the following several categories: (1) Sequence Append: directly appending the visual tokens from different backbones as a longer sequence [20, 33, 55]; (2) Channel Concatenation: concatenating the visual tokens along the channel dimension without increasing the sequence length [49]; (3) LLaVA-HR: injecting high-resolution features into low-resolution vision encoders using mixture-of-resolution adapter [59]; (4) Mini-Gemini: using the CLIP tokens as the low resolution queries to cross-attend another high-resolution vision encoder in the co-located local windows [45]. (5) Deformable Attention: a new baseline we introduce on top of Mini-Gemini, where the vanilla window attention is replaced with deformable attention [86]. Fig. 2 gives a detailed illustration of these fusion strategies. To better study them, we choose “CLIP+ConvNeXt” and “CLIP+ConvNeXt+SAM” as the base multi-encoder combinations to perform comparisons.
Our study in Table 4 shows that Channel Concatenation achieves the best average performance while maintaining better throughput compared to sequence append. The “injection-based” methods, such as LLaVA-HR, Mini-Gemini and Deformable Attention, are in general less competitive on TextVQA [69] and OCRBench [57], performing worse than using ConvNeXt alone as the vision encoder. A plausible explanation is that the CLIP features continue to play a dominant role in the visual tokens. Although sequence append shows comparable performance to channel concatenation, it faces the challenge to handle more vision encoders due to the increasing sequence length. Hence, we choose direct channel concatenation as our fusion strategy considering its performance, expandability, and efficiency.
Table 4: Comparison of different fusion methods for different vision experts. SA, $C$, LH, MG and DA denote the fusion strategies: Sequence Append, Channel Concatenation, LLaVA-HR [59], Mini-Gemini [45] and Deformable Attention, respectively. “Token (V)” denotes the number of vision feature tokens fed into LLM. “#Tokens/s” denotes the speed of the whole pipeline. “Avg” denotes average. Note that Channel Concatenation achieves both the best performance and better throughput.
Vision Encoder(s) Fusion Token (V) #Tokens/s  GQA  MME MMMU  OCR  SQA POPE TextVQA   Avg
CLIP 1024 47.9 64.6 1531 35.7 496 72.8 87.9 64.0 644.4
ConvNeXt 1024 46.8 63.7 1433 37.0 527 72.0 87.9 71.7 652.3
CLIP + ConvNeXt SA 2048 46.1 64.6 1482 35.3 536 72.5 88.3 71.8 657.2
$C$ 1024 47.3 64.0 1486 36.0 533 72.7 88.6 71.9 658.2
LH 1024 47.0 62.9 1488 36.8 521 72.6 88.9 68.7 652.1
MG 1024 45.2 64.9 1481 36.7 474 72.8 88.2 68.6 646.7
DA 1024 47.3 63.6 1497 35.3 459 71.9 88.0 69.1 640.9
CLIP + ConvNeXt + SAM SA 3072 40.3 64.9 1439 35.0 504 72.7 88.2 71.3 649.2
$C$ 1024 46.3 64.6 1497 35.2 558 71.7 87.9 72.2 660.4
2.5 Vison-language Pre-Alignment
Figure 3: The proposed training strategy of Eagle. It consists of three progressive stages, including vision-language pre-alignment training, joint-project training and supervised fine-tuning. These stages effectively leverage public data from diverse sources, ranging from noisy image-text pairs on the web to high-quality caption, VQA, and multimodal dialogue datasets.
As shown in Table 3, vision backbones pre-trained exclusively on vision tasks (e.g., detection, OCR, and segmentation) are less competitive compared to encoder pre-trained on vision language alignment. This is possibly due to representational inconsistencies when integrated with large language models. Additionally, when combining different encoders, there is a gap between these encoders, creating difficulties in the training process. To address this feature inconsistency, we propose a Pre-Alignment training stage that first aligns each individual vision encoder with the same large language model, fostering better synergy between visual and linguistic capabilities.
Fig. 3 depicts our pre-alignment strategy. Instead of training a projector to simultaneously align multiple vision experts as in LLaVA’s [54] original pre-training strategy, we first align the representation of each individual expert with a smaller language model (Vicuna-7B in practice) using next-token-prediction supervision. As shown in Fig. 3, with pre-alignment, the whole training process consists of three steps: 1) training each pre-trained vision expert with their own projector on SFT data, while keeping the language model frozen; 2) combining all the vision experts from the first step and training only the projector with image-text pairs data; 3) training the whole model on the SFT data.
To verify the proposed method, we compare the pre-alignment strategy with the normal two-stage training strategy in Table 5, considering both freezing and unfreezing vision experts for comparison. As shown in Table 5, although unfreezing the vision experts during SFT helps improve performance by updating the vision experts to fit the language model, the Pre-Align strategy more effectively mitigates the inherent biases of each vision expert and stabilizes the training process, subsequently improving overall performance.
Table 5: Experiments to show the effectiveness of Pre-Align. “OCR” denotes OCRBench [57].“Avg” denotes average.
CLIP Vision Expert (X) Unfreeze Pre-Align  MME MMMU  OCR  SQA POPE TextVQA   Avg
CLIP-448 SAM-1024 1495 35.1 292 71.8 85.6 59.7 593
1504 36.9 482 73.5 87.7 63.2 641
1554 36.7 505 73.8 88.2 62.9 649
CLIP-448 ConvNext-1024 1499 35.1 416 72.3 86.7 67.1 629
1522 36.6 550 72.7 87.9 71.9 667
1538 36.3 556 71.9 87.9 71.6 667
CLIP-448 Pix2Struct-1024 1500 35.8 467 71.9 86.1 60.9 627
1529 36.2 535 73.1 87.1 64.5 651
1564 35.3 541 73.3 86.6 63.9 652
CLIP-448 EVA-02-L-1024 1484 35.8 305 72.0 86.0 59.9 597
1537 37.0 474 72.6 87.5 62.6 639
1586 37.3 495 73.1 88.9 64.2 653
2.6 Extension to multi-experts
With the optimized strategies and training recipes of incorporating individual vision experts, we consider the incorporation of even more vision experts to push the limit. To conduct the searching in a systematic and principled manner, we adopt a step-by-step greedy strategy to incorporate additional vision experts.
We consider the vision experts discussed in Section 2.3 for experiments. We mark CLIP, ConvNeXt, SAM, DINOv2, Pix2Struct, and EVA-02-L as A, B, C, D, E, and F, respectively. A round-robin scheme, as shown in Table 6, is adopted. We first use the two top performing vision encoders, CLIP and ConvNeXt, as the basis and gradually add one more vision encoder each time. In each round, the best-performing vision encoder combination is retained for the next round.
Table 6: Results of vision expert selection process. A, B, C, D, E and F denote CLIP, ConvNeXt, SAM, DINOv2, Pix2Struct and EVA-02-L models respectively. “OCR” denotes OCRBench [57]. “Avg” denotes average. Under the same number of encoders, the best scores of various combination variants are bold.
#Encoders Encoder Combination  GQA  MME MMMU  OCR  SQA POPE TextVQA   Avg
2 A + B 64.0 1486 36.0 533 72.7 88.6 71.9 658.3
3 A + B + C 64.6 1497 35.2 558 71.7 87.9 72.2 660.3
A + B + D 65.7 1506 35.0 509 72.6 87.9 70.1 653.6
A + B + E 65.5 1459 35.3 565 73.1 87.9 70.8 660.1
A + B + F 64.7 1506 35.9 562 73.2 88.3 72.1 665.3
4 A + B + F + C 65.9 1498 36.6 540 72.0 88.8 71.6 662.6
A + B + F + D 65.6 1521 35.8 508 73.0 89.0 70.0 657.5
A + B + F + E 65.3 1539 37.1 552 73.0 87.4 71.3 666.1
5 A + B + F + E + C 65.7 1511 37.6 548 72.7 88.6 71.6 666.5
A + B + F + E + D 65.6 1512 36.0 528 72.9 88.4 70.5 660.0
6 A + B + F + E + C + D 65.18 1504 35.3 534 71.9 87.6 70.86 656.3
Table 6 reveals several insights. Generally, introducing additional vision encoders enhances the performance. This indicates that the distinct advantages of different encoders can be preserved and utilized; for example, integrating the EVA-02 encoder improves metrics on the POPE benchmark. Although individual metrics may vary, the aggregated performance shows an upward trend, as evidenced by normalized average metrics, suggesting that the overall efficacy of the system is enhanced with more encoders. Also, Table 6 shows that the best combination of vision experts are CLIP, ConvNeXt, SAM, Pix2Struct, and EVA-02. We will use this recipe in our final model.
Table 7: The advanced training data used in the supervised fine-tuning stage.
Note Data Source
Total Number (1,809k) LLaVA-1.5 [52] (665k), DocVQA [61] (39k), synDog-EN [35] (50k), ChartQA [60] (28k), DVQA [32] (25k), AI2D [34] (15k), ShareGPT-4V [8] (100k), laion-GPT4V [1] (11k), LVIS-Instruct4V [74] (220k), LRV-Instruct [51] (150k), Geo170k [25] (120k), LLaVAR [84] (20k), Visual7W [87] (70k), Open-Hermes 2.5 [72] (1809k)
3 Experiments
After meticulously developing our recipes, we have formulated the following principals for our model, Eagle: (1) integrating more vision experts with optimized training recipe; (2) combining multiple vision experts through direct channel concatenation; (3) pre-training the vision experts separately via pre-alignment. In this section, to further show the advantages of the Eagle models, we add more training data and compare Eagle against the current state-of-the-art MLLMs on different tasks.
3.1 Implementation details
Models. We use Vicuna-v1.5-7B [15], Llama3-8B [3] and Vicuna-v1.5-13B [15] as the language model. For vision encoder, following the results in Sec. 2.6, we denote the model with four vision encoders: CLIP, ConvNeXt, Pix2Struct and EVA-02 as Eagle-X4 and the model with an additional SAM vision encoder as Eagle-X5.
Augmented training data. Considering the training cost, we only use 934K data during the investigation and ablations, as listed in Sec. 2.1. In this section, inspired by recent practices [73, 45], we introduce more datasets, including LVIS-INSTRUCT4V [74], LRV [51], Geo170K [25], LLaVAR [84], Visual7w [87], and OpenHermes-2.5 [72], resulting in a total of 1,809k data for the supervised fine-tuning stage. See Table 7 for more details on this data mixture. To better perform an apple-to-apple comparison with recent open-source MLLM projects, we also train Eagle-X5 with the same data as Cambrian-1 [73], including 2.5 million and 7 million data for pre-training and supervised fine-tuning.
3.2 Main results
Evaluation on visual question answering tasks. We compare Eagle model series across three Visual Question Answering (VQA) benchmarks, including GQA [30], VQAv2 [26] and VizWiz [27]. As shown in Table 8, Eagle-X5 achieves state-of-the-art performance on GQA and VQAv2, underscoring the advantages of incorporating additional vision experts.
Table 8: Results with the advanced training data recipe. SQAI denotes ScienceQA-IMG [65].
Model
MME
MMBench
SEED
MathVista
MMMU
POPE
SQAI
GQA
VizWiz
VQAv2
OCRBench
TextVQA
ChartQA
Vicuna-7B & Qwen-7B LLaVA-1.5 [52] 1510 64.3 58.6 - - 85.9 66.8 62.0* 50.0 78.5* 297 58.2 -
LLaVA-NeXt [53] 1519 67.4 70.2 34.6 35.8 86.5 70.1 64.2* 57.6 80.0* 490 64.9 -
InternVL [13] 1525 - 65.4 - - 86.4 - 62.9* 52.5 79.3* - 57.0 -
LLaVA-HR [59] 1554 - 64.2 - - 87.6 65.1 64.2* 48.7 81.9* - 67.1 -
Monkey [47] - - - - - - - 60.7* 61.2* 80.3* 514 67.6 65.1
Mini-Gemini [45] 1523 65.8 - 32.2 36.8 - 71.1 64.5* - - 477 65.2 -
Eagle-X4 1561 67.8 73.4 37.3 34.9 88.4 70.4 64.8* 50.8 83.4* 540 70.9 67.5
+Pre-Align 1561 68.1 73.4 38.8 36.8 88.7 70.8 64.9* 51.0 83.7* 544 71.9 67.9
Eagle-X5 1528 68.4 73.9 37.0 36.3 88.8 70.0 64.9* 54.4 83.4* 529 71.2 67.7
+Pre-Align 1579 68.8 73.5 39.0 36.4 88.5 71.6 64.0* 56.0 83.6* 554 71.9 67.0
Vicuna-13B LLaVA-1.5 [52] 1531 67.7 61.6 - 36.4 85.9 71.6 63.3* 53.6 80.0* 331 61.3 -
LLaVA-NeXt [53] 1575 70.0 71.9 35.3 36.2 86.2 73.5 65.4* 60.5 82.8* 514 67.1 62.2
InternVL [13] 1546 - - - - 87.1 - 63.9* 54.6 80.2* 517 58.7 -
LLaVA-UHD [79] 1535 68.0 - - - 89.1 72.0 65.2* 56.1 81.7* - 67.7 -
LLaVA-HR [59] 1540 - 64.5 - - 87.8 68.1 64.8* 57.9 82.6* - 68.1 -
Mini-Gemini [45] 1565 68.6 70.6 37.0 37.3 - 71.9 65.8* - - 466 65.9 56.6
Eagle-X4 1627 69.9 74.4 37.6 36.9 87.7 73.1 66.3* 56.2 83.7* 561 73.9 70.5
+Pre-Align 1621 69.9 74.9 39.4 37.3 89.1 73.5 64.9* 55.9 84.5* 581 74.3 72.2
Eagle-X5 1609 69.2 74.1 38.8 36.6 87.8 72.7 66.2* 59.3 83.8* 574 74.2 69.9
+Pre-Align 1604 70.5 74.9 39.7 38.0 88.2 73.1 64.4* 60.9 84.5* 573 73.9 71.0
Evaluation on OCR and chart understanding tasks. To evaluate the OCR, document and chart understanding capabilities of Eagle, we benchmark our model on OCRBench [57], TextVQA [69], and ChartQA [60]. As illustrated in Table 8, our model significantly surpasses competitors on TextVQA, benefiting from its high-resolution architecture and integration of different vision encoders. Notably, Eagle maintains a straightforward design, supporting up to 1024 tokens without requiring complex tile decomposition of images.
Fig. 5 shows some examples of OCR and document understanding cases. With high-resolution adaptation and more vision experts, our model can identifies small text within images and accurately extracts information according to the users’ instructions. To better understand the benefits of introducing experts pre-trained on other vision tasks,we visualize the results of a model with only the ConvNeXt and CLIP vision encoders, compared to the results of Eagle-X5 in Fig. 4. With the full set of vision encoders, the model can successfully correct mistakes, showing that even when equipped with high-resolution vision encoders pre-trained on vision-language alignment, the model’s abilities can still be enhanced by integrating additional vision experts pre-trained on diverse vision tasks.
Evaluation on multimodal benchmarks. We evaluate Eagle on seven benchmarks for MLLMs to demonstrate its capabilities from different perspectives, including MME [24], MMBench [56], SEED [41], MathVista [58], MMMU [81], ScienceQA [65], and POPE [46]. Specifically, MME, MMBench, and SEED assess the overall performance on various real-world tasks based on reasoning, recognition, knowledge, and OCR. MMMU focuses on challenging problems from diverse domains that require college-level knowledge. POPE evaluates the visual hallucinations of MLLMs. The metrics used in our paper adhere to the default settings of these benchmarks. We report the perception score for MME, the en_dev split for MMBench, the image split of SEED, the test-mini split of MathVista, the val split of MMMU, the F1-score of POPE, and the image of SQA to align with the reported scores from other models.
From the data presented in Table 8, Eagle consistently surpasses existing models across various MLLMs on SEED and MME, demonstrating the comprehensive knowledge and reasoning abilities of Eagle. With the help of vision encoders on object-centric tasks, Eagle also achieves the best performance on the POPE benchmark. Additionally, the Pre-Alignment strategy discussed in Sec. 2.5 has been found to further enhance performance when integrating multiple task-specific vision backbones. This approach not only mitigates the inherent biases of each vision expert and the synergy between different modalities but also establishes a robust framework for the multiple-expert fusion.
Comparison with Cambrian-1. Using the same pre-training and supervised fine-tuning datasets from Cambrian-1 [73], Eagle demonstrates superior performance across all the evaluated benchmarks without bells and whistles. As shown in Table 9, Eagle outperforms the Cambrian-1 counterparts considerably on the OCR and Chart category. Consistent improvements are also observed on the General, Knowledge and Vision-Centric categories, showing the robustness and generalization ability of the improved perception design in Eagle.
Table 9: Results using Cambrian-1 training data. The Eagle-X5 models are trained using the same pre-training and supervised fine-tuning data as Cambrian-1 [73]. SQAI denotes ScienceQA-IMG [65]. RWQA denotes RealworldQA [78].
Model Knowledge General OCR and Chart Vision-Centric
Avg
SQAI
MMMU
MathVista
AI2D
Avg
MME
MMBench
SEED
GQA
Avg
ChartQA
OCRBench
TextVQA
DocVQA
Avg
MMVP
RWQA
Llama3-8B
MGM-HD 55.7 75.1 37.3 37.0 73.5 72.7 1606 72.7 73.2 64.5 62.9 59.1 47.7 70.2 74.6 40.4 18.7 62.1
Cambrian-1 61.3 80.4 42.7 49.0 73.0 73.1 1547 75.9 74.7 64.6 71.3 73.3 62.4 71.7 77.8 57.6 51.3 64.2
Eagle 64.2 84.3 43.8 52.7 76.1 73.8 1559 75.9 76.3 64.9 76.6 80.1 62.6 77.1 86.6 69.1 71.6 66.5
Vicuna-13B
MGM-HD 54.1 71.9 37.3 37 70.1 70.7 1597 68.6 70.6 63.7 60.8 56.6 46.6 70.2 69.8 38.4 19.3 57.5
Cambrian-1 60.2 79.3 40.0 48.0 73.6 73.7 1610 75.7 74.4 64.3 71.3 73.8 61.9 72.8 76.8 52.2 41.3 63.0
Eagle 63.0 82.0 41.6 54.4 74.0 74.6 1651 75.7 74.8 65.3 75.1 77.6 61.9 75.5 85.4 61.4 58.0 64.8
Yi-34B
MGM-HD 62.4 77.7 48 43.4 80.5 76.2 1659 80.6 75.3 65.8 68.1 67.6 51.8 74.1 78.9 52.3 37.3 67.2
Cambrian-1 67.0 85.6 49.7 53.2 79.7 76.8 1689 81.4 75.3 65.8 71.9 75.6 60.0 76.7 75.5 60.3 52.7 67.8
Eagle 68.6 85.5 53.2 57.9 79.1 76.3 1677 81.0 75.6 64.9 75.4 77.2 62.4 78.8 83.0 68.3 67.0 69.5
Figure 4: Qualitative comparison of different numbers of vision experts. Baseline means Eagle with only CLIP+ConvNext. More Vision Experts denotes the Eagle-X5 model. We highlight a failure case in RED. BLUE indicates the correct answers. With more vision experts, Eagle can more precisely capture the information in the image and provide the correct answer.
Figure 5: Qualitative samples on OCR and document understanding tasks. Eagle is able to extract useful information from small text.
Figure 6: Qualitative results of Eagle on VQA tasks requiring college-level knowledge or deliberate reasoning.
4 Related work
4.1 Multimodal large language models
Our work is related to the general architecture design of multimodal large language models. Besides the line of representative open-source research mentioned in the introduction section, other notable families of MLLMs include, but are not limited to MiniGPT-4 [85, 7], Lynx [82], Otter [40, 39], Qwen-VL [5], CogVLM [75, 29], VILA [48], GPT-4V [2], Gemini [71], and Llama 3.1 [19]. Depending on how vision signals are integrated into the language model, MLLMs can be broadly categorized into “cross-modal attention” ones and “prefix-tuning” ones [80]. The former injects visual information into different layers of LLMs using cross-modal attention [4, 40], whereas the latter views the visual tokens as part of the language token sequence and directly append them with text embeddings [54, 44, 43]. Our model belongs to the prefix-tuning family by following a LLaVA-styled multimodal architecture. Considering that MLLM is a fast growing field, readers are recommended to refer to more detailed studies and surveys such as [80, 42, 23].
4.2 Vision encoder design for MLLMs
Our work is closely related to research focusing on improved vision encoder designs for MLLM. Early works [54, 44] usually adopted vision encoder pre-trained on vision-language alignment tasks such as CLIP [63] and EVA-CLIP [70]. Stronger vision encoders such as SigLIP [83] and InternVL [13] have been proposed to improve vision-language tasks with improved designs, larger model sizes, and better training recipes. Considering that the models are often pre-trained on low-resolution images and lack the ability to encode fine-grained details, higher resolution adaptation is often performed to increase the MLLM input resolution [9, 10, 6, 12].
Besides higher resolution adaptation, models such as LLaVA-NeXT [53], LLaVA-UHD [79], Monkey [47], InternLM-XComposer [17], and InternVL [12] use either tiling or adaptive tiling to handle high resolution input, where input images are divided into lower-resolution patches and process them separately. Although the ability to handle higher resolution is similarly made possible with the introduction of additional vision experts, its spirit is slightly orthogonal to the tiling techniques in terms of introducing the diversity of visual representations. Both techniques are compatible and can be combined together.
Our work is most related to existing models using multiple vision encoders for improved perception. Mini-Gemini [45] and LLaVA-HR [59] propose to fuse high-resolution visual features into the low-resolution visual tokens. Apart from the resolution issue, these pre-trained vision encoders may lack specific abilities such as reading text and localizing objects. Hence, a series of works have integrated vision models pre-trained on various vision tasks to extend the capabilities of vision encoders. For example, Mousi [20], and Brave [33] fuse visual tokens from different vision encoders by concatenating along the channel or token direction. RADIO [64] proposes a multi-teacher distillation method to unify the abilities of different vision encoders into a unified model. MoAI [37], IVE [28], and Prismer [55] further use the output of vision experts, such as OCR, detection, or depth estimation, to supplement more information for MLLM to generate answers. MoVA [88] devises a routing network to assign an optimal vision model conditioned on the given image and instructions.
5 Conclusion
We conduct an in-depth analysis study on the design space for integrating vision encoders for multimodal large language models. Unlike previous works that focus on designing novel fusion paradigms, we find systematic design choice matters and discover a series of useful techniques. Step by step, we optimize the training recipe of individual vision encoders, identify an extendable and efficient fusing method, and gradually combine vision encoders with different domain knowledge. The results show the importance of basic design space. We hope our work can serve as a new basis and bring new inspiration for the vision encoder design for MLLM.

### v2

@@
-Eagle: Exploring The Design Space for Multimodal LLMs with Mixture of Encoders
+Eagle: Exploring The Design Space for Multi- modal LLMs with Mixture of Encoders
 Abstract
@@
 Eagle is a family of multimodal large language models (MLLMs) with a mixture of vision encoders.
-Left: comparisons between Eagle and existing competitive MLLMs with Vicuna-7B [15] and Llama3-8B [3], with Eagle achieving favorable results on all 12 benchmarks.
+Left: comparisons between Eagle and existing competitive MLLMs with Vicuna-13B (Chiang et al., 2023), with Eagle achieving favorable results on all 13 benchmarks.
 Middle: an evolutionary road map of the design space and advanced training recipes leading to consistent and significant improvements.
@@
 The success of large language models (LLMs) has triggered significant interest in enabling their visual perception capability, such that they could see, understand, and reason in the real world.
-At the center of these multimodal large language models (MLLMs) [23] is a typical design where images are converted into a series of visual tokens by the vision encoders and appended with the text embeddings.
-CLIP [63] is often chosen as the vision encoder since its visual representation is align with the text space by pre-training on image-text pairs.
-Depending on the architectures, training recipes and the way how vision tokens are injected into the language model, there exist various notable families of MLLMs such as Flamingo [4], BLIP [44, 43, 16], PaLI [11], PaLM-E [18] and LLaVA [54, 52].
+At the center of these multimodal large language models (MLLMs) (Fei et al., 2024) is a typical design where images are converted into a series of visual tokens by the vision encoders and appended with the text embeddings.
+CLIP (Radford et al., 2021) is often chosen as the vision encoder since its visual representation is aligned with the text space by pre-training on image-text pairs.
+Depending on the architectures, training recipes, and the way how vision tokens are injected into the language model, there exist various notable families of MLLMs such as Flamingo (Alayrac et al., 2022), BLIP (Li et al., 2022; 2023d; Dai et al., 2024), PaLI (Chen et al., 2023e), PaLM-E (Driess et al., 2023) and LLaVA (Liu et al., 2023d; c).
 Most of these works keep relatively low input resolutions due to the limits on pre-trained vision encoders and LLM sequence length.
-Recent studies [47, 53] show that stronger vision encoder design is important for mitigating MLLM hallucinations [50, 77] and improving resolution-sensitive tasks like optical character recognition (OCR).
+Recent studies (Li et al., 2024c; Liu et al., 2024a) show that stronger vision encoder design is important for mitigating MLLM hallucinations (Liu et al., 2023a; Wu et al., 2024) and improving resolution-sensitive tasks like optical character recognition (OCR).
 A constellation of works thus focuses on enhancing the capability of the vision encoder.
-For example, scaling up the pre-training data and parameters of vision encoder [13] or dividing images into low-resolution patches [53, 67].
+For example, scaling up the pre-training data and parameters of vision encoder (Chen et al., 2023f) or dividing images into low-resolution patches (Liu et al., 2024a; Shi et al., 2024).
 However, these approaches usually introduce large training resources.
-An efficient yet powerful strategy is to mixing visual encoders pre-trained with different tasks and input resolutions, either fusing higher resolution encoders with the CLIP encoder [59, 45], sequentially appending features from different encoders [20, 49], or adopting more complex fusion and routing strategies to make the best of different encoders [37, 88].
-Such a “mixture-of-vision-experts” mode is shown to be effective.
-However, a detailed study of its design space with rigorous ablation is still lacking, which motivates us to revisiting this area.
-Questions such as which vision encoder combination to choose, how to fuse different experts together, and how to adjust training strategies with more vision encoders remain to be answered.
+An efficient yet powerful strategy is to mix visual encoders pre-trained with different tasks and input resolutions, either fusing higher resolution encoders with the CLIP encoder (Luo et al., 2024; Li et al., 2024b), sequentially appending features from different encoders (Fan et al., 2024; Lin et al., 2023b; Karamcheti et al., 2024; Tong et al., 2024), or adopting more complex fusion and routing strategies to make the best of different encoders (Lee et al., 2024; Zong et al., 2024).
+In addition, Prismatic VLM (Liu et al., 2024b) incorporates multiple vision encoders with channel concatenation as part of their design space exploration.
+These “mixture-of-vision-experts” strategies are shown to be effective.
+However, a detailed study focusing on their designs is still lacking.
 To address the above questions, our work systematically investigates the mixture-of-vision-encoders design space for improved MLLM perception.
@@
 1, our exploration of the design space consists of the following steps: 1) Benchmarking various vision encoders and searching recipes for higher resolution adaptation; 2) “Apples to apples” comparison between vision encoder fusion strategies; 3) Progressive identification of the optimal combination of multiple vision encoders; 4) Improved vision expert pre-alignment and data mixture.
-Our study covers the performance of vision encoders pre-trained on different tasks and resolutions (e.g., vision-language alignment [31, 14, 63, 66], self-supervised learning [62], detection [22, 21], segmentation [36], and OCR [38]).
+Our study covers the performance of vision encoders pre-trained on different tasks and resolutions (e.g., vision-language alignment (Ilharco et al., 2021; Cherti et al., 2023; Radford et al., 2021; Schuhmann et al., 2022), self-supervised learning (Oquab et al., 2023), detection (Fang et al., 2023b; a), segmentation (Kirillov et al., 2023), and OCR (Lee et al., 2023)).
 We use a round-robin approach to incorporate additional vision experts.
-Starting with the basic CLIP [63] encoder, we add one additional expert at a time with the best improvement in each round.
+Starting with the basic CLIP (Radford et al., 2021) encoder, we add one additional expert each time with the best improvement in each round.
 Our work is not the first one to leverage multiple vision encoders in MLLM. However, the systematic study leads to several interesting new findings under this setting:
   • Unlocking the vision encoders during MLLM training matters.
-This is in sharp contrast to the LLaVA [54, 52] family and many works that consider multiple vision encoders or teachers [49, 55, 20, 33, 64, 37], where freezing the vision encoders has been a common choice.
-  • Some recently proposed fusion strategies [59, 45] do not show significant advantages despite their advanced designs.
+This is in sharp contrast to the LLaVA (Liu et al., 2023d; c) family and many works that consider multiple vision encoders or teachers (Lin et al., 2023b; Liu et al., 2024b; Fan et al., 2024; Kar et al., 2024; Ranzinger et al., 2024; Lee et al., 2024), where freezing the vision encoders has been a common choice.
+  • Some recently proposed fusion strategies (Luo et al., 2024; Li et al., 2024b) do not show significant advantages despite their advanced designs.
 Instead, we find that straightforward channel concatenation stands out as a simple yet competitive fusion strategy, offering the best efficiency and performance.
-  • Incorporating additional vision experts leads to consistent gain, making it a promising path to systematic enhance MLLM perception besides scaling up single encoders.
+  • Incorporating additional vision experts leads to consistent gain, making it a promising path to systematically enhance MLLM perception besides scaling up vision encoders.
 The improvement is particularly pronounced when vision encoders are unlocked.
-  • We propose a pre-alignment stage where non-text-aligned vision experts are individually fine-tuned with a frozen LLM before trained together.
+  • We propose a pre-alignment stage where non-text-aligned vision experts are individually fine-tuned with a frozen LLM before being trained together.
 This stage is found to enhance the MLLM performance significantly under the mixture-of-vision-encoder design.
 We finally conclude our findings into a family of MLLMs termed Eagle.
-Eagle is evaluated on a series of benchmarks, including visual question answering, OCR/document related tasks, and benchmarks tailored for MLLMs.
+Eagle is evaluated on a series of benchmarks, including visual question answering, OCR/document-related tasks, and benchmarks tailored for MLLMs.
 Our model attains state-of-the-art performance across different benchmarks and demonstrates obvious advantages on OCR and document understanding tasks.
-Using the same pre-train and supervised fine-tuning data from Cambrian-1 [73] - a concurrent family of vision-centric MLLMs sharing similar design spirits, Eagle models overall achieve better performance.
+Using the same pre-train and supervised fine-tuning data from Cambrian-1 (Tong et al., 2024) - a concurrent family of vision-centric MLLMs sharing similar design spirits, Eagle models overall achieve better performance.
 We hope that the Eagle can provide a highly performant and easy-to-reproduce MLLM solution to the community.
@@
 As shown in Fig.
-2, we start from extending the basic CLIP encoder [63] to a set of vision experts with different architectures, pre-training tasks and resolutions.
-With these experts, we then compare different fusion architectures and methods, and study how to optimize the pre-training stratgies given more encoders.
-We also give a detailed analysis in how to select the vision encoders to be integrated.
+2, we start by extending the basic CLIP encoder (Radford et al., 2021) to a set of vision experts with different architectures, pre-training tasks, and resolutions.
+With these experts, we then compare different fusion architectures and methods and study how to optimize the pre-training strategies given more encoders.
+We also give a detailed analysis of how to select the vision encoders to be integrated.
 Finally, we put all the findings together and further extend to multiple expert vision encoders with different resolutions and domain knowledge.
 2.1 Base setup
-We adopt LLaVA’s [54] model architecture as the basis, which consists of a large language model, a vision encoder, and a projection layer.
+We adopt LLaVA-1.5’s (Liu et al., 2023c) model architecture as the basis, which consists of a large language model (Vicuna-v1.5 7B (Chiang et al., 2023)), a vision encoder, and a projection layer.
 The projection layer projects the visual embedding from the vision encoder into the text embedding space.
-We mainly focus on the vision encoder design.
-Figure 2: Overview of the Eagle exploration pipeline.
-We investigate the MLLM design space with multiple vision encoders, aiming to identify optimized design choices and improve the MLLM perception.
-We include various types of vision experts pre-trained on different tasks and resolutions.
-We then systematically compare a series of popular fusion strategies under controlled settings, including Sequence Append, Channel Concatenation, LLaVA-HR [59], Mini-Gemini [45], and Deformable Attention [86].
-Finally, the optimal fusion design is further extended to multiple vision encoders to form a strong MLLM perception.
 Base training data.
-We use the same pre-training data as LLaVA-$1.51.5$ [52], which consists of 595595595595k image text pairs.
-For the supervised fine-tuning stage, we collect data from a series of tasks and convert them into multimodal conversations, including: LLaVA-1.5 [52], Laion-GPT4V [1], ShareGPT-4V [8], DocVQA [61], synDog-EN [35], ChartQA [60], DVQA [32], and AI2D [34], resulting in 934k samples.
-Implementation details.
-We first pre-train the model with image-text pairs for one epoch with a batch size of 256, where the whole model is frozen and only the projector layer is updated.
-In the second stage, we further fine-tune the model on the supervised fine-tuning data for one epoch with a batch size of 128.
-For this exploration, we employ Vicuna-7B as the underlying language model.
-The learning rates are set to 1e-3 for the first stage and 2e-5 for the second stage, respectively.
+We adopt the same pre-training data (LLaVA-595k) as LLaVA-1.5 (Liu et al., 2023c) for the first pre-training stage, which consists of 595595595595k image text pairs.
+To fully examine the potential of different vision experts and fusion methods, instead of using the SFT data from LLaVA-1.5 (Liu et al., 2023c), we collect data from a series of tasks and convert them into multimodal conversations for the supervised fine-tuning (SFT) stage, denoted as Eagle1.8M in Table 1.
+Table 1: Composition of the base supervised fine-tuning data (Eagle1.8M).
+Total Data Size Data Source
+1,809k LLaVA-1.5 (665k) (Liu et al., 2023c), DocVQA (39k) (Mathew et al., 2021), synDog-EN (50k) (Kim et al., 2022), ChartQA (28k) (Masry et al., 2022), DVQA (25k) (Kafle et al., 2018), AI2D (15k) (Kembhavi et al., 2016a), ShareGPT-4V (100k) (Chen et al., 2023b), laion-GPT4V (11k) (lai, 2023), LVIS-Instruct4V (220k) (Wang et al., 2023a), LRV-Instruct (150k) (Liu et al., 2023b), Geo170k (120k) (Gao et al., 2023), LLaVAR (20k) (Zhang et al., 2023), Visual7W (70k) (Zhu et al., 2016), Open-Hermes 2.5 (300k) (Teknium, 2023)
+Base training recipes.
+We start from the LLaVA-1.5 (Liu et al., 2023c) recipe where the model is first pre-trained with image-text pairs for one epoch with a batch size of 256.
+The whole model is frozen and only the projector layer is updated in this pre-training stage.
+In the second stage, we further fine-tune the model on the multi-modal conversation data for one epoch with a batch size of 128.
+The learning rates are set to be 1e-3 for the first stage and 2e-5 for the second stage, respectively.
+Base evaluation.
+To conduct a comprehensive comparison of various methods, we adopt 11 distinct benchmarks that span multiple tasks.
+These benchmarks include 1) General VQA tasks: GQA (Hudson & Manning, 2019), VizWiz (Gurari et al., 2018), MME (Fu et al., 2023), SEED (Li et al., 2023c); 2) OCR/document/chart understanding: OCRBench (Liu et al., 2023f), DocVQA (Mathew et al., 2021), ChartQA (Masry et al., 2022); 3) vision-centric tasks: POPE (Li et al., 2023e), RealWorldQA (xAI, 2024); 4) knowledge-based tasks: ScienceQA (Saikh et al., 2022), AI2D (Kembhavi et al., 2016b).
+To obtain an average score, we normalize each benchmark to a total score of 1,000 and then calculate the average score across all benchmarks.
+Figure 2: Overview of the Eagle exploration.
+In this work, we explore the design space of Multi-Modal Large Language Models (MLLMs) with multiple vision encoders, aiming to identify optimized design choices and enhance MLLM perception.
+We collect a range of vision experts and adapt them for integration into MLLMs.
+A systematic comparison of popular fusion paradigms is then conducted under controlled settings.
+After identifying discrepancies between vision experts pre-trained on different tasks, we optimize the pre-training strategy through a pre-alignment stage and use round-robin searching to determine the optimal combination of vision encoders.
 2.2 Stronger CLIP encoder
-We start our exploration with the CLIP model since it has become the main choice for many MLLMs.
-While CLIP models are known to benefit multimodal tasks, their drawbacks have also been well-studied.
-For instance, many existing MLLMs [52] tend to use the pre-trained CLIP resolutions (such as $224\times 224224 × 224$ or $336\times 336336 × 336$) as their input resolutions.
-In these cases, the encoders often fail to capture fine-grained details that are important for resolution-sensitive tasks like OCR and document understanding [47].
-To handle increased input resolution, a common practice is to use tiling where input images are divided into tiles and encoded separately [53, 47].
-Another simpler method is to directly scale up the input resolution and interpolate the position embeddings of the vision transformer model if needed [9, 10, 6].
+We start our exploration by upgrading the vanilla CLIP Radford et al.
+(2021) model since it has become a standard choice for most of the MLLMs Liu et al.
+(2023d; c).
+While CLIP models are known to benefit multimodal tasks via the text-image alignment, they also have inherent drawbacks.
+For instance, many existing MLLMs (Liu et al., 2023c) tend to use the pre-trained CLIP resolutions (such as $224\times 224224 × 224$ or $336\times 336336 × 336$) as their input resolutions.
+In these cases, the encoders often fail to capture fine-grained details that are important for resolution-sensitive tasks like OCR and document understanding (Li et al., 2024c).
+To handle increased input resolution, a common practice is to use tiling where input images are divided into tiles and encoded separately (Liu et al., 2024a; Li et al., 2024c), or just interpolate the position embedding of the vision transformer model to fit high-resolution inputs (Chen et al., 2023c; d; Beyer et al., 2024).
 We compare these two approaches with frozen/unfrozen vision encoders under different resolutions, with the results shown in Table 2.
 Our findings can be summarized as follows:
-  • Unfreezing the CLIP encoder leads to significant improvement when interpolating to a higher MLLM input resolution that is different from the CLIP pre-training resolution.
-There is also no performance degradation when resolutions remain the same.
-  • When the CLIP encoder is frozen, directly adapting it to a higher MLLM input resolution considerably hurts the performance.
-  • Among the compared strategies, directly interpolating to $448\times 448448 × 448$ with an unfrozen CLIP encoder is shown to be both effective and efficient in terms of performance and cost.
-  • The best CLIP encoder gets close to InternVL in performance despite the significantly smaller model size (300M vs. 6B) and less pre-training data.
-It is worth noting that CLIP-448 allows us to match the setting with LLaVA-HR [59] and InternVL [13] where the CLIP encoders are similarly adapted to take $448\times 448448 × 448$ input and output $10241024$ patch tokens [^1] The latest InternVL 1.0 and InternVL 1.2 models also adopt a pixel shuffle module.
-We remove it in our experiment to ensure full alignment.
-See https://github.com/OpenGVLab/InternVL for details..
-In the following investigation, we stick to this simple strategy of scaling up the input resolution and unlocking the vision encoder during training.
-Table 1: Comparison of different high-resolution adaption methods.
-Direct interpolation is more efficient and effective.
-“Res” denotes input image resolution.
-“#Token/s” denotes the inference speed of the whole framework.
-“#Token (V)” denotes the number of visual feature tokens fed into LLM.
-“OCR” denotes OCRBench [57].
-“Avg” denotes average.
-Note that CLIP-448 gets very close to InternVL in terms of performance, despite the significantly smaller model size (300M vs.
-6B).
-Method Unfreeze  Res #Token/s #Token (V)  GQA  MME MMMU  OCR  SQA POPE  Avg
-Original 336 47.2 576 62.46 1488 36.1 369 72.79 86.77 615
-Original 336 47.2 576 63.68 1496 36.6 369 72.22 85.9 617
-Interpolate 448 47.9 1024 62.45 1484 34.4 285 72.27 86.54 598
-Interpolate 448 47.9 1024 64.62 1531 35.7 496 72.77 87.88 645
-Interpolate 672 44.5 2304 64.51 1492 36.2 492 72.86 87.49 641
-Tiled-input 672 44.8 2304 63.59 1455 34.8 445 71.94 87.83 625
-InternVL [13] 448 46.2 1024 65.12 1521 36.2 527 72.46 87.13 649
-Table 2: Detailed setting of the pre-trained vision experts.
-Expert Category Resolution Token Number Post-Processing Variant Hyper-links
-CLIP Image-Text Matching 448 1024 None VIT-L
-ConvNeXt Image Classification 1024 1024 None ConvNeXt-XXL
-EVA-02 Object Detection 1024 1024 Resize EVA-02-Large
-Pix2Struct Text Recognition 1024 1024 Resize Pix2Struct-02-Large
-DINOv2 Self-supervised 448 1024 None dinov2_vitl14_reg
-SAM Image Segmentation 1024 1024 Pixel-unshuffle SAM-Large
+  • Updating the CLIP encoder during SFT significantly improves performance at higher resolutions but slightly reduces it when using the pre-training resolution.
+  • Interpolating CLIP encoder to fit the input size of $448\times 448448 × 448$ offers a strong balance between efficiency and performance, trailing the $672\times 672672 × 672$ version with less than half the tokens.
+  • Despite its smaller size (0.3B vs. 5.9B) and less pre-training data, the CLIP encoder gets close with interpolation approaches InternVL’s (Chen et al., 2023f) performance under the same setting.
+Table 2: Comparing different high-resolution adaption methods.
+“#Tok(V)” denotes the number of visual tokens.
+“#Parames”, “FLOPs” and “Img/Sec” denote the model size, complexity and throughput (bs=4) of the vision encoder.
+Method Unfreeze Res. #Tok(V) #Params FLOPs Img/Sec Avg.
+Original 336 576 0.3B 119G 197.2 616.5
+Original 336 576 0.3B 119G 197.2 562.6
+Interpolate 448 1024 0.3B 214G 119.5 589.7
+Interpolate 448 1024 0.3B 214G 119.5 670.5
+Interpolate 672 2304 0.3B 480G 56.3 674.2
+Tiled-input 672 2304 0.3B 476G 51.6 673.9
+InternVL 448 1024 5.9B 5669G 13.52 661.9
+InternVL 448 1024 5.9B 5669G 13.52 671.5
+Based on the results, we can see that Direct interpolation to $448\times 448448 × 448$ can achieve competitive performance while being more efficient.
+We thus use the CLIP encoder with a $448\times 448448 × 448$ input resolution while unlocking the encoder during the SFT stage.
 2.3 Vision experts
@@
 This also helps us identify the distinct advantages of different experts.
-We collect a set of vision experts, including: (1) Vision-Language Alignment: CLIP [63] and ConvNeXt [76] from OpenCLIP [31, 66].
-(2) Object-Centric: EVA-02 [22, 21] pre-trained on detection datasets.
-(3) OCR: Pix2Struct [38].(4) Segmentation: SAM [36].
-(5) Self-supervised: DINOv2 [62].
-The detailed input resolution and checkpoint of each vision encoder can be found in Table 2.
-We resize the output 2D feature maps of each vision encoder using bilinear interpolation and pixel shuffle [68] to ensure that the visual token number equals $10241024$.
-The results in Table 3 reveal that MLLMs with these task-specific vision encoders achieve optimal performance in their pre-training domains.
-EVA-02 excels in the object hallucination evaluation benchmark POPE [46] and general visual question answering benchmark GQA [30].
+We collect a set of vision experts, including: (1) Vision-Language Alignment: CLIP (Radford et al., 2021) and ConvNeXt (Liu et al., 2022) from OpenCLIP (Ilharco et al., 2021; Schuhmann et al., 2022).
+(2) Object-Centric: EVA-02 (Fang et al., 2023b; a) pre-trained on detection datasets.
+(3) OCR: Pix2Struct (Lee et al., 2023).(4) Segmentation: SAM (Kirillov et al., 2023).
+(5) Self-supervised: DINOv2 (Oquab et al., 2023).
+We resize the output 2D feature maps of each vision encoder using bilinear interpolation or pixel shuffle (Shi et al., 2016) to ensure that the visual token number equals $10241024$.
+Table 3: Comparison between different vision experts as the MLLM encoders.
+Category Vision Encoder   Res. Post-process Unfreeze   Avg. Model Link
+VL Alignment ConvNeXt 1024 None 654.6 ConvNeXt-XXL
+682.1
+Segmentation SAM 1024 Pixel Unshuffle 486.2 SAM-ViT-Large
+510.5
+Object Detection EVA-02 1024 Resize 543.7 EVA-02-L-Det
+639.1
+Text Recognition Pix2Struct 1024 Resize 598.6 Pix2Struct-02-Large
+606.2
+Self-Supervised DINOv2 448 None 520.7 ViT-L/14-Reg
+537.3
+Results in Table 3 show that unfreezing the vision experts again leads to consistent improvement, which is aligned with Sec.
+2.2.
+In addition, results in Table 10 (see Appendix A.1) further demonstrate that MLLMs with these task-specific vision encoders achieve optimal performance in their pre-training domains.
+EVA-02 excels in the object hallucination evaluation benchmark POPE and general visual question answering benchmark GQA.
 CLIP and ConvNeXt perform well across all benchmarks, benefiting from their training on large-scale image-text pairs using contrastive loss.
 Conversely, while Pix2Struct excels in text recognition, it shows limited capability in object recognition and general VQA tasks, like POPE and GQA.
-DINOv2 and SAM, pre-trained with self-supervised learning and semantic segmentation respectively, struggle with text recognition tasks.
-Note that the conclusion made in Sec.
-2.2 aligns with the results here.
-Both ConvNeXt and EVA-02 show significant performance gains when the vision encoders are updated during training.
-Table 3: Comparison between different vision experts as the MLLM encoders. “Res” denotes resolution. “OCR” denotes OCRBench [57].
-Category Vision Tower Unfreeze    Res  GQA  MME MMMU  OCR  SQA POPE TextVQA
-VL Alignment ConvNeXt 1024 63.71 1473 34.2 402 71.92 87.42 66.95
-1024 63.71 1433 37.0 527 72.01 87.91 71.66
-Segmentation SAM 1024 54.87 1193 34.8 32 70.34 83.31 45.04
-1024 60.05 1305 35.9 38 70.64 86.62 45.71
-Object Detection EVA-02 1024 63.54 1422 35.8 58 70.86 88.50 47.86
-1024 64.93 1474 35.7 387 71.78 88.96 59.79
-Text Recognition Pix2Struct 1024 52.47 1217 35.7 443 69.61 77.54 56.56
-1024 54.05 1220 34.3 425 70.29 79.21 57.05
-Self-supervised DINOv2 448 60.81 1269 34.0 39 70.76 86.42 45.71
-448 64.06 1371 34.2 37 71.35 87.29 46.77
+DINOv2 and SAM, pre-trained with self-supervised learning and segmentation, struggle with text recognition tasks.
 2.4 Fusion strategy
-Existing MLLM frameworks have proposed various fusion strategies to ensemble the encoders in Table 3, with the hope that their domain-specific strengths can be leveraged.
+Existing MLLM frameworks have proposed various mixture-of-vision-encoder strategies, with the hope that their domain-specific strengths can be leveraged.
 In all cases, improvements in MLLM performance have been reported with the fusion of vision encoders.
@@
 It is thus not entirely clear how much improvement is from the fusion strategies themselves versus the improved representations from various encoders.
-We notice that existing popular fusion strategies, despite their variations in designs, can be broadly represented by the following several categories: (1) Sequence Append: directly appending the visual tokens from different backbones as a longer sequence [20, 33, 55]; (2) Channel Concatenation: concatenating the visual tokens along the channel dimension without increasing the sequence length [49]; (3) LLaVA-HR: injecting high-resolution features into low-resolution vision encoders using mixture-of-resolution adapter [59]; (4) Mini-Gemini: using the CLIP tokens as the low resolution queries to cross-attend another high-resolution vision encoder in the co-located local windows [45].
-(5) Deformable Attention: a new baseline we introduce on top of Mini-Gemini, where the vanilla window attention is replaced with deformable attention [86].
+We notice that existing popular fusion strategies, despite their variations in designs, can be broadly represented by the following several categories: (1) Sequence Append: directly appending the visual tokens from different backbones as a longer sequence (Fan et al., 2024; Kar et al., 2024); (2) Channel Concatenation: concatenating the visual tokens along the channel dimension without increasing the sequence length (Lin et al., 2023b; Karamcheti et al., 2024); (3) LLaVA-HR: injecting high-resolution features into low-resolution vision encoders using mixture-of-resolution adapter (Luo et al., 2024); (4) Mini-Gemini: using the CLIP tokens as the low-resolution queries to cross-attend another high-resolution vision encoder in the co-located local windows (Li et al., 2024b).
+(5) Deformable Attention: a new baseline we introduce on top of Mini-Gemini, where the vanilla window attention is replaced with deformable attention (Zhu et al., 2021).
 Fig.
@@
 To better study them, we choose “CLIP+ConvNeXt” and “CLIP+ConvNeXt+SAM” as the base multi-encoder combinations to perform comparisons.
-Our study in Table 4 shows that Channel Concatenation achieves the best average performance while maintaining better throughput compared to sequence append.
-The “injection-based” methods, such as LLaVA-HR, Mini-Gemini and Deformable Attention, are in general less competitive on TextVQA [69] and OCRBench [57], performing worse than using ConvNeXt alone as the vision encoder.
-A plausible explanation is that the CLIP features continue to play a dominant role in the visual tokens.
-Although sequence append shows comparable performance to channel concatenation, it faces the challenge to handle more vision encoders due to the increasing sequence length.
-Hence, we choose direct channel concatenation as our fusion strategy considering its performance, expandability, and efficiency.
-Table 4: Comparison of different fusion methods for different vision experts.
-SA, $C$, LH, MG and DA denote the fusion strategies: Sequence Append, Channel Concatenation, LLaVA-HR [59], Mini-Gemini [45] and Deformable Attention, respectively.
-“Token (V)” denotes the number of vision feature tokens fed into LLM.
-“#Tokens/s” denotes the speed of the whole pipeline.
-“Avg” denotes average.
-Note that Channel Concatenation achieves both the best performance and better throughput.
-Vision Encoder(s) Fusion Token (V) #Tokens/s  GQA  MME MMMU  OCR  SQA POPE TextVQA   Avg
-CLIP 1024 47.9 64.6 1531 35.7 496 72.8 87.9 64.0 644.4
-ConvNeXt 1024 46.8 63.7 1433 37.0 527 72.0 87.9 71.7 652.3
-CLIP + ConvNeXt SA 2048 46.1 64.6 1482 35.3 536 72.5 88.3 71.8 657.2
-$C$ 1024 47.3 64.0 1486 36.0 533 72.7 88.6 71.9 658.2
-LH 1024 47.0 62.9 1488 36.8 521 72.6 88.9 68.7 652.1
-MG 1024 45.2 64.9 1481 36.7 474 72.8 88.2 68.6 646.7
-DA 1024 47.3 63.6 1497 35.3 459 71.9 88.0 69.1 640.9
-CLIP + ConvNeXt + SAM SA 3072 40.3 64.9 1439 35.0 504 72.7 88.2 71.3 649.2
-$C$ 1024 46.3 64.6 1497 35.2 558 71.7 87.9 72.2 660.4
+Table 4: Comparison of different fusion methods for different vision experts. “#Token(V)” denotes the number of visual tokens. “#Tokens/s” denotes the inference throughtput of the whole pipeline.
+Vision Encoders Fusion #Token(V) #Tokens/s #Params Avg.
+CLIP + ConvNeXt Seq. Append 2048 46.1 1200M 690.5
+Channel Concat. 1024 47.3 1184M 681.5
+LLaVA-HR 1024 47.0 1219M 678.7
+Mini-Gemini 1024 45.3 1201M 672.5
+Deformable Attn. 1024 47.3 1201M 674.3
+CLIP + ConvNeXt + SAM Seq. Append 3072 40.3 1529M 686.2
+Channel Concat. 1024 46.3 1495M 690.4
+Our study in Table 4 shows that Channel Concatenation stands out with the best performance, expandability, and efficiency.
+The “injection-based” methods, such as LLaVA-HR, Mini-Gemini and Deformable Attention, are in general less competitive on TextVQA (Singh et al., 2019) and OCRBench (Liu et al., 2023f), performing worse than using ConvNeXt alone as the vision encoder.
+Although sequence append shows comparable performance to channel concatenation, it faces the challenge of handling significantly increased sequence lengths with additional vision encoders.
 2.5 Vison-language Pre-Alignment
@@
 These stages effectively leverage public data from diverse sources, ranging from noisy image-text pairs on the web to high-quality caption, VQA, and multimodal dialogue datasets.
-As shown in Table 3, vision backbones pre-trained exclusively on vision tasks (e.g., detection, OCR, and segmentation) are less competitive compared to encoder pre-trained on vision language alignment.
+As shown in Table 3, encoders pre-trained exclusively on vision tasks (e.g., detection, OCR, and segmentation) are less competitive compared to those pre-trained on vision language alignment.
 This is possibly due to representational inconsistencies when integrated with large language models.
@@
 3 depicts our pre-alignment strategy.
-Instead of training a projector to simultaneously align multiple vision experts as in LLaVA’s [54] original pre-training strategy, we first align the representation of each individual expert with a smaller language model (Vicuna-7B in practice) using next-token-prediction supervision.
+Instead of training a projector to simultaneously align multiple vision experts as in LLaVA’s (Liu et al., 2023c) original pre-training strategy, we first align the representation of each individual expert with a smaller language model (Vicuna-7B in practice) using next-token-prediction supervision.
 As shown in Fig.
-3, with pre-alignment, the whole training process consists of three steps: 1) training each pre-trained vision expert with their own projector on SFT data, while keeping the language model frozen; 2) combining all the vision experts from the first step and training only the projector with image-text pairs data; 3) training the whole model on the SFT data.
+3, with pre-alignment, the whole training process consists of three steps: 1) training each pre-trained vision expert with their own projector, while keeping the language model frozen; 2) combining all vision experts from the first step and training both the projector and vision experts; 3) training the whole model on SFT data.
+Table 5: The effectiveness of Pre-alignment.
+CLIP Vision Expert (X) Unfreeze Pre-align Avg.
+CLIP-448 SAM-1024 630.6
+648.5
+662.5
+672.3
+CLIP-448 ConvNext-1024 652.0
+670.1
+681.5
+686.2
+CLIP-448 Pix2Struct-1024 653.5
+665.7
+673.7
+680.4
+CLIP-448 EVA-02-L-1024 630.2
+645.2
+659.2
+668.2
 To verify the proposed method, we compare the pre-alignment strategy with the normal two-stage training strategy in Table 5, considering both freezing and unfreezing vision experts for comparison.
 As shown in Table 5, although unfreezing the vision experts during SFT helps improve performance by updating the vision experts to fit the language model, the Pre-Align strategy more effectively mitigates the inherent biases of each vision expert and stabilizes the training process, subsequently improving overall performance.
-Table 5: Experiments to show the effectiveness of Pre-Align. “OCR” denotes OCRBench [57].“Avg” denotes average.
-CLIP Vision Expert (X) Unfreeze Pre-Align  MME MMMU  OCR  SQA POPE TextVQA   Avg
-CLIP-448 SAM-1024 1495 35.1 292 71.8 85.6 59.7 593
-1504 36.9 482 73.5 87.7 63.2 641
-1554 36.7 505 73.8 88.2 62.9 649
-CLIP-448 ConvNext-1024 1499 35.1 416 72.3 86.7 67.1 629
-1522 36.6 550 72.7 87.9 71.9 667
-1538 36.3 556 71.9 87.9 71.6 667
-CLIP-448 Pix2Struct-1024 1500 35.8 467 71.9 86.1 60.9 627
-1529 36.2 535 73.1 87.1 64.5 651
-1564 35.3 541 73.3 86.6 63.9 652
-CLIP-448 EVA-02-L-1024 1484 35.8 305 72.0 86.0 59.9 597
-1537 37.0 474 72.6 87.5 62.6 639
-1586 37.3 495 73.1 88.9 64.2 653
 2.6 Extension to multi-experts
 With the optimized strategies and training recipes of incorporating individual vision experts, we consider the incorporation of even more vision experts to push the limit.
-To conduct the searching in a systematic and principled manner, we adopt a step-by-step greedy strategy to incorporate additional vision experts.
+To conduct the search in a systematic manner, we adopt a step-by-step greedy strategy to incorporate additional vision experts.
 We consider the vision experts discussed in Section 2.3 for experiments.
-We mark CLIP, ConvNeXt, SAM, DINOv2, Pix2Struct, and EVA-02-L as A, B, C, D, E, and F, respectively.
-A round-robin scheme, as shown in Table 6, is adopted.
-We first use the two top performing vision encoders, CLIP and ConvNeXt, as the basis and gradually add one more vision encoder each time.
+We mark CLIP, ConvNeXt, SAM, DINOv2, Pix2Struct, and EVA-02-L as CL, CN, SA, DI, PS, and EV, respectively.
+A round-robin scheme, as shown in Fig.
+4, is adopted.
+We first use the two top-performing vision encoders, CLIP and ConvNeXt, as the basis and gradually add one more vision encoder each time.
 In each round, the best-performing vision encoder combination is retained for the next round.
-Table 6: Results of vision expert selection process.
-A, B, C, D, E and F denote CLIP, ConvNeXt, SAM, DINOv2, Pix2Struct and EVA-02-L models respectively.
-“OCR” denotes OCRBench [57].
-“Avg” denotes average.
-Under the same number of encoders, the best scores of various combination variants are bold.
-#Encoders Encoder Combination  GQA  MME MMMU  OCR  SQA POPE TextVQA   Avg
-2 A + B 64.0 1486 36.0 533 72.7 88.6 71.9 658.3
-3 A + B + C 64.6 1497 35.2 558 71.7 87.9 72.2 660.3
-A + B + D 65.7 1506 35.0 509 72.6 87.9 70.1 653.6
-A + B + E 65.5 1459 35.3 565 73.1 87.9 70.8 660.1
-A + B + F 64.7 1506 35.9 562 73.2 88.3 72.1 665.3
-4 A + B + F + C 65.9 1498 36.6 540 72.0 88.8 71.6 662.6
-A + B + F + D 65.6 1521 35.8 508 73.0 89.0 70.0 657.5
-A + B + F + E 65.3 1539 37.1 552 73.0 87.4 71.3 666.1
-5 A + B + F + E + C 65.7 1511 37.6 548 72.7 88.6 71.6 666.5
-A + B + F + E + D 65.6 1512 36.0 528 72.9 88.4 70.5 660.0
-6 A + B + F + E + C + D 65.18 1504 35.3 534 71.9 87.6 70.86 656.3
-Table 6 reveals several insights.
+#Encoder Encoder Combination Config #Params (M) FLOPs (G) Img/Sec Avg.
+2 CL + CN X2 1155.2 3347.2 18.1 681.5
+3 CL + CN + DI 1460.6 3659.9 15.1 685.4
+CL + CN + SA 1463.9 4657.8 8.8 690.4
+CL + CN + PS 1669.6 4373.2 6.9 685.1
+CL + CN + EV X3 1459.6 4280.9 9.1 690.7
+4 CL + CN + EV + DI 1765.1 4593.6 8.3 688.0
+CL + CN + EV + SA 1768.4 5591.5 5.9 689.4
+CL + CN + EV + PS X4 1974.1 5306.9 5.0 694.6
+5 CL + CN + EV + PS + DI 2279.5 5619.5 4.7 684.7
+CL + CN + EV + PS + SA X5 2282.8 6617.4 3.8 697.1
+6 CL + CN + EV + PS + SA + DI X6 2588.2 6930.1 3.6 686.8
+Figure 4: Results of vision expert selection process.
+CL, CN, EV, PS, SA and DI denote CLIP, ConvNeXt, EVA-02, Pix2Struct, SAM and DINOv2, respectively.
+(Left) The performance of various vision encoder combinations is presented, highlighting how different pairings influence overall effectiveness.
+“#Parames”, “FLOPs” and “Img/Sec” denote the model size, complexity and throughput (bs=4) of the vision encoder.
+(Right) The curve illustrates the average score as the number of vision encoders increases.
+Each point on the curve represents the best-performing combination for the corresponding number of vision encoders.
+No Pre-Alignment is used in this comparison.
+Fig.
+4 reveals several insights.
 Generally, introducing additional vision encoders enhances the performance.
@@
 Although individual metrics may vary, the aggregated performance shows an upward trend, as evidenced by normalized average metrics, suggesting that the overall efficacy of the system is enhanced with more encoders.
-Also, Table 6 shows that the best combination of vision experts are CLIP, ConvNeXt, SAM, Pix2Struct, and EVA-02.
-We will use this recipe in our final model.
-Table 7: The advanced training data used in the supervised fine-tuning stage.
-Note Data Source
-Total Number (1,809k) LLaVA-1.5 [52] (665k), DocVQA [61] (39k), synDog-EN [35] (50k), ChartQA [60] (28k), DVQA [32] (25k), AI2D [34] (15k), ShareGPT-4V [8] (100k), laion-GPT4V [1] (11k), LVIS-Instruct4V [74] (220k), LRV-Instruct [51] (150k), Geo170k [25] (120k), LLaVAR [84] (20k), Visual7W [87] (70k), Open-Hermes 2.5 [72] (1809k)
+Also, Fig.
+4 shows that the best combination of vision experts are CLIP, ConvNeXt, SAM, Pix2Struct, and EVA-02.
+We will use this recipe in the final model.
 3 Experiments
-After meticulously developing our recipes, we have formulated the following principals for our model, Eagle: (1) integrating more vision experts with optimized training recipe; (2) combining multiple vision experts through direct channel concatenation; (3) pre-training the vision experts separately via pre-alignment.
-In this section, to further show the advantages of the Eagle models, we add more training data and compare Eagle against the current state-of-the-art MLLMs on different tasks.
+In this section, we take the findings and the best-explored design from Section 2 and compare them against the current state-of-the-art MLLMs on different tasks.
 3.1 Implementation details
-Models.
-We use Vicuna-v1.5-7B [15], Llama3-8B [3] and Vicuna-v1.5-13B [15] as the language model.
-For vision encoder, following the results in Sec.
-2.6, we denote the model with four vision encoders: CLIP, ConvNeXt, Pix2Struct and EVA-02 as Eagle-X4 and the model with an additional SAM vision encoder as Eagle-X5.
-Augmented training data.
-Considering the training cost, we only use 934K data during the investigation and ablations, as listed in Sec.
-2.1.
-In this section, inspired by recent practices [73, 45], we introduce more datasets, including LVIS-INSTRUCT4V [74], LRV [51], Geo170K [25], LLaVAR [84], Visual7w [87], and OpenHermes-2.5 [72], resulting in a total of 1,809k data for the supervised fine-tuning stage.
-See Table 7 for more details on this data mixture.
-To better perform an apple-to-apple comparison with recent open-source MLLM projects, we also train Eagle-X5 with the same data as Cambrian-1 [73], including 2.5 million and 7 million data for pre-training and supervised fine-tuning.
+Language models. We use Vicuna-v1.5-7B (Chiang et al., 2023), Llama3-8B (AI@Meta, 2024) and Vicuna-v1.5-13B (Chiang et al., 2023) as the language models.
+Vision encoders.
+We follow the best X4 and X5 configurations, where the interpolated CLIP-448 and pre-aligned vision experts are channel-concatenated, and trained following the exact best training recipes in Figure 3 and Table 5.
+Training recipe. On Eagle1.8M, we follow the base recipe in Section 2.1 with encoder learning rate the same as SFT (2e-5). On Cambrian data, we follow Tong et al. (2024) with PT/SFT bs=1024.
 3.2 Main results
 Evaluation on visual question answering tasks.
-We compare Eagle model series across three Visual Question Answering (VQA) benchmarks, including GQA [30], VQAv2 [26] and VizWiz [27].
-As shown in Table 8, Eagle-X5 achieves state-of-the-art performance on GQA and VQAv2, underscoring the advantages of incorporating additional vision experts.
-Table 8: Results with the advanced training data recipe. SQAI denotes ScienceQA-IMG [65].
+We compare Eagle model series across three Visual Question Answering (VQA) benchmarks, including GQA (Hudson & Manning, 2019), VQAv2 (Goyal et al., 2017) and VizWiz (Gurari et al., 2018).
+As shown in Table 6, Eagle-X5 achieves state-of-the-art performance on GQA and VQAv2, underscoring the advantages with additional vision experts.
+Table 6: Main results with base training data. SQAI denotes image split of ScienceQA.
 Model
 MME
-MMBench
+MMB
 SEED
@@
 VQAv2
-OCRBench
+OCR
 TextVQA
 ChartQA
-Vicuna-7B & Qwen-7B LLaVA-1.5 [52] 1510 64.3 58.6 - - 85.9 66.8 62.0* 50.0 78.5* 297 58.2 -
-LLaVA-NeXt [53] 1519 67.4 70.2 34.6 35.8 86.5 70.1 64.2* 57.6 80.0* 490 64.9 -
-InternVL [13] 1525 - 65.4 - - 86.4 - 62.9* 52.5 79.3* - 57.0 -
-LLaVA-HR [59] 1554 - 64.2 - - 87.6 65.1 64.2* 48.7 81.9* - 67.1 -
-Monkey [47] - - - - - - - 60.7* 61.2* 80.3* 514 67.6 65.1
-Mini-Gemini [45] 1523 65.8 - 32.2 36.8 - 71.1 64.5* - - 477 65.2 -
-Eagle-X4 1561 67.8 73.4 37.3 34.9 88.4 70.4 64.8* 50.8 83.4* 540 70.9 67.5
-+Pre-Align 1561 68.1 73.4 38.8 36.8 88.7 70.8 64.9* 51.0 83.7* 544 71.9 67.9
+Vicuna-7B & Qwen-7B LLaVA-1.5 (Liu et al., 2023c) 1510 64.3 58.6 - - 85.9 66.8 62.0* 50.0 78.5* 297 58.2 -
+LLaVA-NeXt (Liu et al., 2024a) 1519 67.4 70.2 34.6 35.8 86.5 70.1 64.2* 57.6 80.0* 490 64.9 -
+InternVL (Chen et al., 2023f) 1525 - 65.4 - - 86.4 - 62.9* 52.5 79.3* - 57.0 -
+LLaVA-HR (Luo et al., 2024) 1554 - 64.2 - - 87.6 65.1 64.2* 48.7 81.9* - 67.1 -
+Monkey (Li et al., 2024c) - - - - - - - 60.7* 61.2* 80.3* 514 67.6 65.1
+Mini-Gemini (Li et al., 2024b) 1523 65.8 - 32.2 36.8 - 71.1 64.5* - - 477 65.2 -
 Eagle-X5 1528 68.4 73.9 37.0 36.3 88.8 70.0 64.9* 54.4 83.4* 529 71.2 67.7
-+Pre-Align 1579 68.8 73.5 39.0 36.4 88.5 71.6 64.0* 56.0 83.6* 554 71.9 67.0
-Vicuna-13B LLaVA-1.5 [52] 1531 67.7 61.6 - 36.4 85.9 71.6 63.3* 53.6 80.0* 331 61.3 -
-LLaVA-NeXt [53] 1575 70.0 71.9 35.3 36.2 86.2 73.5 65.4* 60.5 82.8* 514 67.1 62.2
-InternVL [13] 1546 - - - - 87.1 - 63.9* 54.6 80.2* 517 58.7 -
-LLaVA-UHD [79] 1535 68.0 - - - 89.1 72.0 65.2* 56.1 81.7* - 67.7 -
-LLaVA-HR [59] 1540 - 64.5 - - 87.8 68.1 64.8* 57.9 82.6* - 68.1 -
-Mini-Gemini [45] 1565 68.6 70.6 37.0 37.3 - 71.9 65.8* - - 466 65.9 56.6
-Eagle-X4 1627 69.9 74.4 37.6 36.9 87.7 73.1 66.3* 56.2 83.7* 561 73.9 70.5
-+Pre-Align 1621 69.9 74.9 39.4 37.3 89.1 73.5 64.9* 55.9 84.5* 581 74.3 72.2
+Eagle-X5 (+Pre-Align) 1582 69.7 73.7 38.2 38.0 88.7 71.9 64.6* 58.7 83.6* 566 71.9 69.3
+Vicuna-13B LLaVA-1.5 (Liu et al., 2023c) 1531 67.7 61.6 - 36.4 85.9 71.6 63.3* 53.6 80.0* 331 61.3 -
+LLaVA-NeXt (Liu et al., 2024a) 1575 70.0 71.9 35.3 36.2 86.2 73.5 65.4* 60.5 82.8* 514 67.1 62.2
+InternVL (Chen et al., 2023f) 1546 - - - - 87.1 - 63.9* 54.6 80.2* 517 58.7 -
+LLaVA-UHD (Xu et al., 2024) 1535 68.0 - - - 89.1 72.0 65.2* 56.1 81.7* - 67.7 -
+LLaVA-HR (Luo et al., 2024) 1540 - 64.5 - - 87.8 68.1 64.8* 57.9 82.6* - 68.1 -
+Mini-Gemini (Li et al., 2024b) 1565 68.6 70.6 37.0 37.3 - 71.9 65.8* - - 466 65.9 56.6
 Eagle-X5 1609 69.2 74.1 38.8 36.6 87.8 72.7 66.2* 59.3 83.8* 574 74.2 69.9
-+Pre-Align 1604 70.5 74.9 39.7 38.0 88.2 73.1 64.4* 60.9 84.5* 573 73.9 71.0
+Eagle-X5 (+Pre-Align) 1605 71.6 74.9 42.7 38.5 89.2 75.5 64.6* 60.9 84.5* 598 73.3 72.1
 Evaluation on OCR and chart understanding tasks.
-To evaluate the OCR, document and chart understanding capabilities of Eagle, we benchmark our model on OCRBench [57], TextVQA [69], and ChartQA [60].
+To evaluate the OCR, document, and chart understanding capabilities of Eagle, we benchmark our model on OCRBench (Liu et al., 2023f), TextVQA (Singh et al., 2019), and ChartQA (Masry et al., 2022).
 As illustrated in Table 6, our model significantly surpasses competitors on TextVQA, benefiting from its high-resolution architecture and integration of different vision encoders.
-Notably, Eagle maintains a straightforward design, supporting up to 1024 tokens without requiring complex tile decomposition of images.
+Notably, Eagle maintains a straightforward design, supporting up to 1024x1024 resolution without requiring complex tile decomposition of images.
 Fig.
 5 shows some examples of OCR and document understanding cases.
-With high-resolution adaptation and more vision experts, our model can identifies small text within images and accurately extracts information according to the users’ instructions.
+With high-resolution adaptation and more vision experts, our model can identify small text within images and accurately extract information according to the users’ instructions.
 To better understand the benefits of introducing experts pre-trained on other vision tasks, we visualize the results of a model with only the ConvNeXt and CLIP vision encoders, compared to the results of Eagle-X5 in Fig.
-4.
+5.
 With the full set of vision encoders, the model can successfully correct mistakes, showing that even when equipped with high-resolution vision encoders pre-trained on vision-language alignment, the model’s abilities can still be enhanced by integrating additional vision experts pre-trained on diverse vision tasks.
 Evaluation on multimodal benchmarks.
-We evaluate Eagle on seven benchmarks for MLLMs to demonstrate its capabilities from different perspectives, including MME [24], MMBench [56], SEED [41], MathVista [58], MMMU [81], ScienceQA [65], and POPE [46].
+We evaluate Eagle on seven benchmarks for MLLMs to demonstrate its capabilities from different perspectives, including MME (Fu et al., 2023), MMBench (Liu et al., 2023e), SEED (Li et al., 2023c), MathVista (Lu et al., 2024), MMMU (Yue et al., 2024), ScienceQA (Saikh et al., 2022), and POPE (Li et al., 2023e).
 Specifically, MME, MMBench, and SEED assess the overall performance on various real-world tasks based on reasoning, recognition, knowledge, and OCR.
@@
 The metrics used in our paper adhere to the default settings of these benchmarks.
-We report the perception score for MME, the en_dev split for MMBench, the image split of SEED, the test-mini split of MathVista, the val split of MMMU, the F1-score of POPE, and the image of SQA to align with the reported scores from other models.
+We report the perception score for MME, the en_dev split for MMBench, the image split of SEED, the test-mini split of MathVista, the val split of MMMU, the F1-score of POPE, and the image split of SQA to align with the reported scores from other models.
+Table 7: Comparison between different training strategies.
+“1 epoch” means we train Eagle for 1 epoch in the supervised fine-tuning stage.
+‘unlock*” means we unlock vision encoders in the pre-training stage.
+More details in Table.
+13.
+Config Summary Pre-align Pre-train Fine-tune Avg.
+1 epoch LLaVA-595K Eagle1.8M 697.1
+2 epoch LLaVA-595K Eagle1.8M 698.3
+1 epoch, unlock* LLaVA-595K Eagle1.8M 698.0
+1 epoch, unlock* LLaVA-595K+Eagle1.8M Eagle1.8M 699.5
+1 epoch Eagle1.8M LLaVA-595K Eagle1.8M 706.6
+1 epoch, unlock* Eagle1.8M LLaVA-595K Eagle1.8M 707.1
+1 epoch, unlock* LLaVA-595K+Eagle1.8M LLaVA-595K Eagle1.8M 707.8
+1 epoch, unlock* LLaVA-595K+Eagle1.8M LLaVA-595K+Eagle1.8M Eagle1.8M 708.9
 From the data presented in Table 6, Eagle consistently surpasses existing models across various MLLMs on SEED and MME, demonstrating the comprehensive knowledge and reasoning abilities of Eagle.
@@
 2.5 has been found to further enhance performance when integrating multiple task-specific vision backbones.
-This approach not only mitigates the inherent biases of each vision expert and the synergy between different modalities but also establishes a robust framework for the multiple-expert fusion.
+This approach not only mitigates the inherent biases of each vision expert and the synergy between different modalities but also establishes a robust framework for multiple-expert fusion.
+Study on more advanced training recipes.
+Table.
+7 presents our step-by-step experiments to study the training recipes.
+We found that the best recipe is to first pre-align each vision expert on LLaVA-595K + Eagle1.8M.
+In the pretraining stage, we combine all vision experts from the first step and training both the projector and vision experts on LLaVA-595K + Eagle1.8M.
+Finally, we train the whole model on the Eagle1.8M.
 Comparison with Cambrian-1.
-Using the same pre-training and supervised fine-tuning datasets from Cambrian-1 [73], Eagle demonstrates superior performance across all the evaluated benchmarks without bells and whistles.
-As shown in Table 9, Eagle outperforms the Cambrian-1 counterparts considerably on the OCR and Chart category.
-Consistent improvements are also observed on the General, Knowledge and Vision-Centric categories, showing the robustness and generalization ability of the improved perception design in Eagle.
-Table 9: Results using Cambrian-1 training data.
-The Eagle-X5 models are trained using the same pre-training and supervised fine-tuning data as Cambrian-1 [73].
-SQAI denotes ScienceQA-IMG [65].
-RWQA denotes RealworldQA [78].
+Using the same pre-training and supervised fine-tuning datasets from Cambrian-1 (Tong et al., 2024), Eagle demonstrates superior performance across all the evaluated benchmarks without bells and whistles.
+As shown in Table 8, Eagle outperforms the Cambrian-1 counterparts considerably for the OCR and Chart category.
+Consistent improvements are also observed for the General, Knowledge, and Vision-Centric categories, showing the robustness and generalization ability of the improved perception design in Eagle.
+Table 8: Results using the same training data as Cambrian-1 (Tong et al., 2024). SQAI denotes ScienceQA-IMG (Saikh et al., 2022). RWQA denotes the RealworldQA (xAI, 2024).
 Model Knowledge General OCR and Chart Vision-Centric
@@
 MME
-MMBench
+MMB
 SEED
@@
 ChartQA
-OCRBench
+OCR
 TextVQA
@@
 Cambrian-1 61.3 80.4 42.7 49.0 73.0 73.1 1547 75.9 74.7 64.6 71.3 73.3 62.4 71.7 77.8 57.6 51.3 64.2
-Eagle 64.2 84.3 43.8 52.7 76.1 73.8 1559 75.9 76.3 64.9 76.6 80.1 62.6 77.1 86.6 69.1 71.6 66.5
+Eagle-X5 65.2 84.1 43.5 56.9 76.2 74.0 1587 75.5 76.5 64.9 77.0 80.7 62.6 76.7 87.1 59.6 52.0 67.2
 Vicuna-13B
-MGM-HD 54.1 71.9 37.3 37 70.1 70.7 1597 68.6 70.6 63.7 60.8 56.6 46.6 70.2 69.8 38.4 19.3 57.5
+MGM-HD 54.1 71.9 37.3 37.0 70.1 70.7 1597 68.6 70.6 63.7 60.8 56.6 46.6 70.2 69.8 38.4 19.3 57.5
 Cambrian-1 60.2 79.3 40.0 48.0 73.6 73.7 1610 75.7 74.4 64.3 71.3 73.8 61.9 72.8 76.8 52.2 41.3 63.0
-Eagle 63.0 82.0 41.6 54.4 74.0 74.6 1651 75.7 74.8 65.3 75.1 77.6 61.9 75.5 85.4 61.4 58.0 64.8
+Eagle-X5 63.8 82.6 42.2 54.6 73.8 74.6 1651 75.7 75.0 65.0 75.7 78.6 62.4 74.9 86.7 54.8 44.6 65.0
 Yi-34B
-MGM-HD 62.4 77.7 48 43.4 80.5 76.2 1659 80.6 75.3 65.8 68.1 67.6 51.8 74.1 78.9 52.3 37.3 67.2
+MGM-HD 62.4 77.7 48.0 43.4 80.5 76.2 1659 80.6 75.3 65.8 68.1 67.6 51.8 74.1 78.9 52.3 37.3 67.2
 Cambrian-1 67.0 85.6 49.7 53.2 79.7 76.8 1689 81.4 75.3 65.8 71.9 75.6 60.0 76.7 75.5 60.3 52.7 67.8
-Eagle 68.6 85.5 53.2 57.9 79.1 76.3 1677 81.0 75.6 64.9 75.4 77.2 62.4 78.8 83.0 68.3 67.0 69.5
+Eagle-X5 68.6 85.5 53.2 57.9 79.1 76.3 1677 81.0 75.6 64.9 75.4 77.2 62.4 78.8 83.0 59.8 50.0 69.5
+4 Related work
+Multimodal large language models.
+Our work is related to the general architecture design of multimodal large language models.
+Besides the line of representative open-source research mentioned in the introduction section, other notable families of MLLMs include, but are not limited to MiniGPT-4 (Zhu et al., 2023; Chen et al., 2023a), Lynx (Zeng et al., 2023), Otter (Li et al., 2023b; a), Qwen-VL (Bai et al., 2023), CogVLM (Wang et al., 2023b; Hong et al., 2024), VILA (Lin et al., 2023a), GPT-4V (Achiam et al., 2023), Gemini (Team et al., 2023), and Llama 3.1 (Dubey et al., 2024).
+Depending on how vision signals are integrated into the language model, MLLMs can be broadly categorized into “cross-modal attention” ones and “prefix-tuning” ones (Yin et al., 2024).
+The former injects visual information into different layers of LLMs using cross-modal attention (Alayrac et al., 2022; Li et al., 2023b), whereas the latter views the visual tokens as part of the language token sequence and directly append them with text embeddings (Liu et al., 2023d; Li et al., 2022; 2023d).
+Our model belongs to the prefix-tuning family by following a LLaVA-styled multimodal architecture.
+Considering that MLLM is a fast-growing field, readers are recommended to refer to more detailed studies and surveys such as Yin et al.
+(2024); Li et al.
+(2024a); Fei et al.
+(2024).
+Vision encoder design for MLLMs.
+Our work is closely related to research focusing on improved vision encoder designs for MLLM.
+Early works (Liu et al., 2023d; Li et al., 2022) usually adopted vision encoder pre-trained on vision-language alignment tasks such as CLIP (Radford et al., 2021) and EVA-CLIP (Sun et al., 2023).
+Stronger vision encoders such as SigLIP (Zhai et al., 2023) and InternVL (Chen et al., 2023f) have been proposed to improve vision-language tasks with improved designs, larger model sizes, and better training recipes.
+Considering that the models are often pre-trained on low-resolution images and lack the ability to encode fine-grained details, higher resolution adaptation is often performed to increase the MLLM input resolution (Chen et al., 2023c; d; Beyer et al., 2024; Chen et al., 2024).
+Besides higher resolution adaptation, models such as LLaVA-NeXT (Liu et al., 2024a), LLaVA-UHD (Xu et al., 2024), Monkey (Li et al., 2024c), InternLM-XComposer (Dong et al., 2024), and InternVL (Chen et al., 2024) use adaptive tiling to divide high-resolution input into lower resolution patches that are processed separately.
+Although the ability to handle higher resolution is similarly made possible with the introduction of additional vision experts, its spirit is slightly orthogonal to the tiling techniques in terms of diversifying the visual representations.
+Both techniques are compatible and can be combined together.
+Our work is most related to existing models using multiple vision encoders for improved perception.
+Mini-Gemini (Li et al., 2024b) and LLaVA-HR (Luo et al., 2024) propose to fuse high-resolution visual features into the low-resolution visual tokens.
+Apart from the resolution issue, these pre-trained vision encoders may lack specific abilities such as reading text and localizing objects.
+Hence, a series of works have integrated vision models pre-trained on different tasks for more comprehensive capabilities.
+For example, Mousi (Fan et al., 2024), Prismatic VLM (Karamcheti et al., 2024) and Brave (Kar et al., 2024) fuse visual tokens from different vision encoders by concatenating along the channel or token direction.
+There are also more complex approaches, including knowledge distillation Ranzinger et al.
+(2024), augmenting the input prompt with the information output by the vision experts (Lee et al., 2024; He et al., 2024; Liu et al., 2024b), or using a routing network to assign input to proper vision experts (Zong et al., 2024).
+In particular, Prismatic VLM (Karamcheti et al., 2024) systematically explores the design space of MLLMs across various dimensions, including data, training recipes, and notably.
+It also includes the vision encoder ensemble as part of its enhancement strategies.
+However, it lacks a comprehensive ablation study and discussion addressing the challenges associated with combining multiple vision encoders.
 Figure 5: Qualitative comparison of different numbers of vision experts.
@@
 With more vision experts, Eagle can more precisely capture the information in the image and provide the correct answer.
-Figure 5: Qualitative samples on OCR and document understanding tasks. Eagle is able to extract useful information from small text.
-Figure 6: Qualitative results of Eagle on VQA tasks requiring college-level knowledge or deliberate reasoning.
-4 Related work
-4.1 Multimodal large language models
-Our work is related to the general architecture design of multimodal large language models.
-Besides the line of representative open-source research mentioned in the introduction section, other notable families of MLLMs include, but are not limited to MiniGPT-4 [85, 7], Lynx [82], Otter [40, 39], Qwen-VL [5], CogVLM [75, 29], VILA [48], GPT-4V [2], Gemini [71], and Llama 3.1 [19].
-Depending on how vision signals are integrated into the language model, MLLMs can be broadly categorized into “cross-modal attention” ones and “prefix-tuning” ones [80].
-The former injects visual information into different layers of LLMs using cross-modal attention [4, 40], whereas the latter views the visual tokens as part of the language token sequence and directly append them with text embeddings [54, 44, 43].
-Our model belongs to the prefix-tuning family by following a LLaVA-styled multimodal architecture.
-Considering that MLLM is a fast growing field, readers are recommended to refer to more detailed studies and surveys such as [80, 42, 23].
-4.2 Vision encoder design for MLLMs
-Our work is closely related to research focusing on improved vision encoder designs for MLLM.
-Early works [54, 44] usually adopted vision encoder pre-trained on vision-language alignment tasks such as CLIP [63] and EVA-CLIP [70].
-Stronger vision encoders such as SigLIP [83] and InternVL [13] have been proposed to improve vision-language tasks with improved designs, larger model sizes, and better training recipes.
-Considering that the models are often pre-trained on low-resolution images and lack the ability to encode fine-grained details, higher resolution adaptation is often performed to increase the MLLM input resolution [9, 10, 6, 12].
-Besides higher resolution adaptation, models such as LLaVA-NeXT [53], LLaVA-UHD [79], Monkey [47], InternLM-XComposer [17], and InternVL [12] use either tiling or adaptive tiling to handle high resolution input, where input images are divided into lower-resolution patches and process them separately.
-Although the ability to handle higher resolution is similarly made possible with the introduction of additional vision experts, its spirit is slightly orthogonal to the tiling techniques in terms of introducing the diversity of visual representations.
-Both techniques are compatible and can be combined together.
-Our work is most related to existing models using multiple vision encoders for improved perception.
-Mini-Gemini [45] and LLaVA-HR [59] propose to fuse high-resolution visual features into the low-resolution visual tokens.
-Apart from the resolution issue, these pre-trained vision encoders may lack specific abilities such as reading text and localizing objects.
-Hence, a series of works have integrated vision models pre-trained on various vision tasks to extend the capabilities of vision encoders.
-For example, Mousi [20], and Brave [33] fuse visual tokens from different vision encoders by concatenating along the channel or token direction.
-RADIO [64] proposes a multi-teacher distillation method to unify the abilities of different vision encoders into a unified model.
-MoAI [37], IVE [28], and Prismer [55] further use the output of vision experts, such as OCR, detection, or depth estimation, to supplement more information for MLLM to generate answers.
-MoVA [88] devises a routing network to assign an optimal vision model conditioned on the given image and instructions.
 5 Conclusion
@@
 Unlike previous works that focus on designing novel fusion paradigms, we find systematic design choice matters and discover a series of useful techniques.
-Step by step, we optimize the training recipe of individual vision encoders, identify an extendable and efficient fusing method, and gradually combine vision encoders with different domain knowledge.
+Step by step, we optimize the training recipes of individual vision encoders, identify an extendable and efficient fusion method, and gradually combine vision encoders with different domain knowledge.
 The results show the importance of basic design space.
 We hope our work can serve as a new basis and bring new inspiration for the vision encoder design for MLLM.
+6 Acknowledgments
+The team would like to thank Yunhao Fang and Jason Lu for sharing the data and training recipes, and Yawen Luo for the assistance on figure editing.
+We thank Wei Ping, Zhuolin Yang, Wenliang Dai, Nayeon Lee, Boxin Wang, Ilia Karmanov, Lukas Voegtle, Philipp Fischer, Matthieu Le and Tuomas Rintamaki for their assistance on the internal codebases.
+We also thank the valuable discussions and input from Zhiqi Li, Guo Chen, Shilong Liu, Jihao Liu, Ming-Chang Chiu, Yunze Man, Shiyi Lan, Nadine Chang, Maying Shen, Vibashan VS, Jenny Schmalfuss, Jose Alvarez, Amala Sanjay Deshmukh, Mike Ranzinger, Greg Heinrich, Pavlo Molchanov, Vidya Murali, Parthasarathy Sriram, Mohammad Shoeybi, Song Han, Ofri Masad, Osvald Nitski, Qing Miao, Yao Xu, Jane Scowcroft, Dmitry Chichkov and Padmavathy Subramanian.
+Finally, the team would like to thank the Hugging Face Team that for the support of ZERO GPU demo, and the NVIDIA infrastructure team for their prompt and helpful assistance.
+Min Shi is partly supported by NSF Award #2427478 and #2229873.
+Appendix A Appendix
+A.1 Benchmark details
+In this section, we provide the additional benchmark details for the tables in Section 2.
+The detailed comparison of different adaptation methods for the CLIP encoder are shown in Table 9.
+Table 10 shows the comparison between different vision encoders on all the adopted benchmarks.
+Table 11 list the detailed results on the vision encoder fusion methods.
+Table 12 shows the comparison between different vision encoder combinations on each benchmark.
+Table 9: Comparison of different high-resolution adaption methods to strengthen CLIP model (336x336). RWQA denotes the RealworldQA (xAI, 2024).
+Vision Encoder Unfreeze Res GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
+Original 336 63.2 55.1 1574 70.2 354 59.6 42.1 71.1 86.7 58.0 72.5 616.5
+Original 336 60.9 54.0 1501 61.8 305 56.4 24.5 69.3 80.9 53.6 72.0 562.6
+Interpolate 448 62.7 53.8 1470 69.1 299 58.4 37.7 71.2 85.7 55.3 69.6 589.7
+Interpolate 448 65.6 57.8 1534 73.7 526 65.0 61.1 73.7 87.3 57.7 71.5 670.5
+Interpolate 672 64.9 55.7 1503 72.4 509 64.6 62.0 72.2 87.1 57.4 71.2 674.2
+Tiled-input 672 63.0 54.9 1529 72.5 435 64.9 65.7 71.5 87.6 57.0 71.4 673.9
+InternVL 448 63.6 56.9 1537 71.7 529 65.0 58.6 72.9 87.4 59.2 70.2 661.9
+InternVL 448 65.6 57.8 1534 73.7 526 65.0 61.1 73.7 87.3 58.8 71.5 671.5
+Table 10: Comparison between different vision experts as the MLLM encoders.
+Category Vision Encoder Unfreeze Res GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
+VL Alignment ConvNeXt 1024 63.3 53.5 1526 70.6 404 70.4 60.8 71.6 87.5 57.1 68.6 635.0
+1024 63.3 54.4 1510 72.9 518 77.9 67.0 72.1 88.1 58.8 68.6 659.7
+Segmentation SAM 1024 57.3 49.0 1216 56.9 38 20.1 17.4 69.2 84.3 49.2 66.8 471.1
+1024 60.2 51.5 1291 65.9 35 21.2 17.8 70.7 86.4 54.1 65.7 494.7
+Object Detection EVA-02 1024 63.1 51.1 1359 69.2 123 25.6 25.2 71.2 88.5 57.9 66.1 523.5
+1024 64.3 55.5 1449 72.7 358 57.1 57.5 72.2 88.3 59.6 67.7 614.4
+Text Recognition Pix2Struct 1024 53.1 48.1 1296 53.4 460 71.0 61.0 69.6 79.2 46.7 65.5 578.7
+1024 54.9 47.3 1262 55.1 472 72.5 62.0 68.7 80.0 49.3 66.6 584.8
+Self-Supervised DINOv2 448 62.4 53.1 1438 67.4 41 20.2 17.3 70.7 85.3 53.3 67.1 503.1
+448 64.2 55.5 1466 71.8 45 20.4 17.5 71.4 87.4 57.3 67.8 518.0
+Table 11: Comparison of different fusion methods for different vision experts. “#Tokens(V)” denotes the number of visual tokens.
+Vision Encoders Fusion #Tokens(V) GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
+CLIP + ConvNeXt Seq. Append 2048 64.8 54.5 1563 73.4 532 77.7 67.6 72.4 87.9 61.2 68.8 690.5
+Channel Concat. 1024 63.2 48.0 1497 73.5 551 77.7 67.0 72.4 88.3 59.1 70.7 681.5
+LLaVA-HR 1024 64.5 57.2 1538 72.0 498 74.5 63.8 72.3 87.7 59.2 68.7 678.7
+Mini-Gemini 1024 65.3 56.9 1548 72.9 478 68.3 63.2 71.5 87.3 59.7 69.4 672.5
+Deformable Attn. 1024 64.0 57.3 1504 72.7 463 69.5 64.4 73.3 87.4 62.8 68.9 674.3
+CLIP + ConvNeXt + SAM Seq. Append 3072 64.3 53.6 1539 73.2 525 77.9 67.0 72.3 87.4 60.1 69.5 686.2
+Channel Concat. 1024 63.3 55.9 1528 73.3 545 78.9 67.2 72.3 88.4 59.2 70.0 690.4
+Table 12: Detailed comparison on vision encoder combinations.
+#Encoder Encoder Combination GQA VizWiz MME SEED OCR DocVQA ChartQA AI2D POPE RWQA SQA Avg.
+2 CL + CN 63.2 48.0 1497.0 73.5 551.0 77.7 67.0 72.4 88.3 59.1 70.7 681.5
+3 CL + CN + DI 63.3 55.9 1528.0 73.3 545.0 78.9 67.2 72.3 88.4 59.2 70.0 690.4
+CL + CN + SA 64.6 55.3 1504.0 73.3 526.0 75.7 64.9 72.1 88.3 61.1 70.9 685.4
+CL + CN + PS 63.2 51.4 1497.0 73.3 550.0 78.5 65.9 73.1 87.7 60.3 70.5 685.1
+CL + CN + EV 63.2 51.7 1565.0 73.9 538.0 77.7 67.8 73.6 89.0 61.4 69.4 690.7
+CL + CN + EV + DI 63.6 54.9 1512.0 73.8 547.0 77.0 66.7 73.1 88.9 60.4 69.7 689.4
+CL + CN + EV + SA 64.3 57.7 1533.0 73.7 521.0 75.2 65.3 72.2 88.5 61.1 70.2 688.0
+CL + CN + EV + PS 64.8 56.5 1561.0 73.4 540.0 78.8 67.5 72.2 88.4 59.9 70.5 694.6
+5 CL + CN + EV + PS + SA 64.7 59.1 1528.0 73.9 529.0 78.6 67.8 72.9 88.8 62.2 69.7 697.1
+CL + CN + EV + PS + DI 64.7 54.1 1506.0 73.7 541.0 75.1 64.9 72.7 88.3 60.0 70.3 684.7
+6 CL + CN + EV + PS + SA + DI 63.8 57.8 1512.0 73.5 525.0 75.1 65.8 71.8 88.4 61.4 69.9 686.8
+Table 13: Comparison between different training strategies.
+‘1 epoch” means we train Eagle for 1 epoch in the supervised fine-tuning stage.
+‘unlock*” means we unlock vision encoders in the pre-training stage.
+Config Prealign Pretrain Finetune
+GQA
+MME
+OCR
+SciQA
+POPE
+DocVQA
+ChartQA
+SEED
+Vizwiz
+AI2D
+RWQA
+Avg.
+1 epoch llava595k Eagle1.8M 64.7 1528 52.9 69.7 88.8 78.6 67.7 73.9 59.1 72.8 62.2 697.1
+2 epoch llava595k Eagle1.8M 65.4 1539 51.4 70.3 87.9 79.8 67.9 73.8 58.5 73.5 62.7 698.3
+1 epoch, unlock* llava595k Eagle1.8M 64.1 1541 54.4 71.5 88.5 79.1 68.5 74.0 56.6 72.2 61.9 698.0
+1 epoch, unlock* llava595k+Eagle1.8M Eagle1.8M 65.3 1545 54.8 70.5 88.5 78.8 68.4 73.5 57.5 72.0 62.9 699.5
+1 epoch Eagle1.8M llava595k Eagle1.8M 64.5 1579 55.4 71.6 88.7 79.3 68.9 73.9 58.9 73.9 63.4 706.6
+1 epoch, unlock* Eagle1.8M llava595k Eagle1.8M 65.1 1561 56.0 71.3 88.5 79.3 69.5 74.2 59 73.5 63.4 707.1
+1 epoch, unlock* llava595k+Eagle1.8M llava595k Eagle1.8M 64.5 1570 56.0 71.7 89 79.7 68.9 74.0 58.5 74.0 63.7 707.8
+1 epoch, unlock* llava595k+Eagle1.8M llava595k+Eagle1.8M Eagle1.8M 64.6 1582 56.6 71.9 88.7 79.8 69.3 73.7 58.7 73.5 63.9 708.9
+A.2 Additional qualitative results
+Fig. 6 demonstrates that our model can perform reasoning based on the knowledge to answer some difficult questions. Fig. 7 shows some examples of OCR and document understanding cases.
+Figure 6: Qualitative results of Eagle on VQA tasks requiring college-level knowledge or deliberate reasoning.
+Figure 7: Qualitative samples on OCR and document understanding tasks. Eagle is able to extract useful information from small text.